### v1

# Stability of heat kernel bounds under pointed Gromov-Hausdorff convergence

Abstract
We present a method for constructing a heat kernel on the pointed Gromov-Hausdorff limit space and demonstrate the stability of heat kernel estimates under this convergence. Furthermore, we establish the Mosco convergence of the associated energy forms along a subsequence.
1 Introduction
Heat kernels, also known as transition densities, play a central role in the analysis of manifolds, metric spaces, and broader mathematical structures. In diverse geometric contexts, such as Riemannian manifolds, Alexandrov spaces, and metric measure spaces, heat kernels give valuable information about the geometric and analytical properties of the underlying space. Understanding their behavior becomes particularly relevant when studying the convergence of a sequence of these spaces.
The convergence of spaces can occur in multiple frameworks, such as the construction of fractals or the study of the tangent and asymptotic cones of general metric measure spaces, which can be formulated in the language of Gromov-Hausdorff topology. In [KK94, KK96, Din02, Xu14], the convergence of heat kernels, Green functions, and the Harnack inequality for Riemannian manifolds under Gromov-Hausdorff convergence is studied. In [CHK17], it is shown that if a sequence of measurable spaces equipped with resistance metric satisfies uniform volume doubling (equivalent to certain heat kernel bounds when the resistance forms are local, as shown in [Kum04]), and also Gromov-Hausdorff converges, then the corresponding stochastic processes converge in distribution. As a result, the same type of heat kernel estimates holds for the limiting process.
In this work, we focus on the convergence of the following type of heat kernel bounds in the context of strongly local Dirichlet forms on metric measure spaces. For detailed definitions and notation on Dirichlet forms, see Section 4 or [FOT11].
Definition 1.1.
Let $(X,d,m)$ be a metric measure space and $\Psi:(0,\infty)\to(0,\infty)$ be a continuous increasing bijection of $(0,\infty)$ onto itself. We say that a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X,m)$ satisfies the heat kernel estimates $\operatorname{HKE}(\Psi)$, if there exists $C_{1},c_{2},c_{3},\delta\in(0,\infty)$ and a heat kernel $\left\{p_{t}\right\}_{t>0}$ of its semigroup $\{P_{t}\}_{t>0}$ such that for any $t>0$,
$p_{t}(x,y)$ $\leq\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-c_{2}t\Phi\left(c_{3}\frac{d(x,y)}{t}\right)\right)\qquad\mbox{for $m$-a.e.\ $x,y\in X$}$ (1.1)
and $\text{and }\ p_{t}(x,y)$ $\geq\frac{C_{1}^{-1}}{V(x,\Psi^{-1}(t))}\qquad\mbox{for $m$-a.e.\ $x,y\in X$ with $d(x,y)\leq\delta\Psi^{-1}(t)$},$ (1.2)
where $V(x,r):=m(B(x,r))$ and
$\Phi(s):=\Phi_{\Psi}(s):=\sup_{r>0}\left({\frac{s}{r}-\frac{1}{\Psi(r)}}\right).$ (1.3)
The heat kernel estimate $\mathrm{HKE}(\Psi_{\beta})$, where $\Psi_{\beta}(r)=r^{\beta}$ arises in various contexts. For $\beta=2$, it corresponds to the Gaussian heat kernel estimate, which holds for Riemannian manifolds with non-negative Ricci curvature [LY86]. For $\beta>2$, it corresponds to the sub-Gaussian heat kernel estimate, which is valid for diffusions on many fractals, including the Sierpiński gasket [BP88] and the Sierpiński carpet [BB92].
The main result of this paper is as follows. See Sections 2 and 7 for the definitions of pointed Gromov-Hausdorff convergence and Mosco convergence, respectively.
Theorem 1.2.
[…]
  1. (A1)
    The Dirichlet form $(\mathcal{E}_{j},\mathcal{F}_{j})$ satisfies $\operatorname{HKE}(\Psi_{j})$ for each $1\leq j<\infty$, with uniform constant $C_{1}$, $C_{2}$ and $\delta$.
  2. (A2)
    For all $1\leq j\leq\infty$ and all $0<r\leq R<\infty$,
    $C^{-1}_{0}\left({\frac{R}{r}}\right)^{\beta}\leq\frac{\Psi_{j}(R)}{\Psi_{j}(r)}\leq C_{0}\left({\frac{R}{r}}\right)^{\beta^{\prime}}$ (1.4)
    for some constants $1<\beta\leq\beta^{\prime}$, $C_{0}>1$.
  3. (A3)
    $\Psi_{j}\to\Psi_{\infty}$ uniformly on compact set on $[0,\infty)$.
  4. (A4)
    There exists $\{V_{j}:(0,\infty)\to(0,\infty):1\leq j\leq\infty\}$, $V_{l}:(0,\infty)\to(0,\infty)$ and $V_{u}:(0,\infty)\to(0,\infty)$ that are increasing functions, and there exist constants $\alpha,\alpha^{\prime}>0$, $C_{\mathrm{v}}\geq 1$ such that
    $C_{\mathrm{v}}^{-1}\left({\frac{R}{r}}\right)^{\alpha}\leq\frac{V_{j}(R)}{V_{j}(r)}\leq C_{\mathrm{v}}\left({\frac{R}{r}}\right)^{\alpha^{\prime}},\quad\forall 0<r\leq R<\infty,j\in\mathbb{N}\cup\{u,l\},$ (1.5)
[…]
    and
[…]
  5. (A5)
    $V_{j}\to V_{\infty}$ uniformly on compact set on $[0,\infty)$.
  6. (A6)
    Assume further that $(X_{j},d_{j})$, $1\leq j<\infty$ are proper length spaces, and for some points $p_{j}\in X_{j}$, $1\leq j<\infty$, there holds
[…]
    for some complete metric space $(X_{\infty},d_{\infty})$ and some $p_{\infty}\in X_{\infty}$.
Then the following holds:
  1. (1)
    There exists a Radon measure $m_{\infty}$ on the Borel sigma algebra of $(X_{\infty},d_{\infty})$ such that for some $C\geq 1$,
[…]
  2. (2)
    There exists a conservative, regular and strongly local symmetric Dirichlet form $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ on $L^{2}(X_{\infty},m_{\infty})$ satisfying $\operatorname{HKE}(\Psi_{\infty})$.
  3. (3)
    Along a subsequence, $\mathcal{E}_{j}\rightarrow\mathcal{E}_{\infty}$ with respect to the Mosco topology, as $j\to\infty$.
Remark 1.3.
  1. (i)
    If a metric measure space $(X,d,m)$ satisfies the chain condition [GT12, Definition 6.9] (for example, if $(X,d)$ is a proper length space [KM20, Proposition A.1]), and a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X,m)$ satisfies $\operatorname{HKE}(\Psi)$, then the heat kernel estimates can be improved to the full heat kernel bound $\mathrm{HKE}_{\mathrm{f}}(\Psi)$ [GT12, Theorem 6.5]. Specifically, the upper bound of $p_{t}$ in $\operatorname{HKE}(\Psi)$ remains valid, while the lower bound of $p_{t}$ in $\operatorname{HKE}(\Psi)$ is replaced by
    $p_{t}(x,y)\geq\frac{C_{1}^{-1}}{V(x,\Psi^{-1}(t))}\exp\left(-c_{4}t\Phi\left(c_{5}\frac{d(x,y)}{t}\right)\right)\qquad\mbox{for $m$-a.e.\ $x,y\in X$}$
    for some $C_{1},c_{4},c_{5}\in(0,\infty)$.
    By [HKST15, Proposition 11.3.12, Proposition 11.3.14], (A6) implies that $(X_{\infty},d_{\infty})$ is also a proper length space. Therefore, the full heat kernel bound $\mathrm{HKE}_{\mathrm{f}}(\Psi_{j})$ holds for $(\mathcal{E}_{j},\mathcal{F}_{j})$ for all $1\leq j\leq\infty$.
  2. (ii)
    Since the assumptions and conclusions in Theorem 1.2 are stable under a bi-Lipschitz change of metric, we may instead assume that each $d_{j}$ is (uniformly) bi-Lipschitz equivalent to a length metric, which can be ensured by quasi-convexity [HKST15, Proposition 8.3.12] or by the chain condition [KM20, Proposition A.1].
  3. (iii)
    It is possible to prove similar results by assuming beforehand that the existence of a measure $m_{\infty}$ on $X_{\infty}$ and that the sequence $(X_{n},d_{n},p_{n},m_{n})$ pointed measured Gromov-Hausdorff converges111see [HKST15, Definition 11.4.5] or [KS03, Definition 2.2] for definitions. to $(X_{\infty},d_{\infty},p_{\infty},m_{\infty})$. However, it appears that directly verifying pointed measured Gromov-Hausdorff convergence is more challenging.
We briefly outline the proof of Theorem 1.2. The Radon measure $m_{\infty}$ on limit space is derived from the tightness of the push-forward measures. To obtain the limit Dirichlet form $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$, we first expand the heat kernel of the part of $(\mathcal{E}_{n},\mathcal{F}_{n})$ on the balls $B_{n}(p_{n},R)$ with Dirichlet boundary conditions, using their eigenvalues and eigenfunctions (see (4.1)). We then apply a version of Arzelà–Ascoli theorem (Lemma 5.1) to take the limit and obtain the limits of eigenvalues and eigenfunctions, which gives a heat kernel on $B_{\infty}(p_{\infty},R)$. Letting $R\to\infty$, we obtain a heat kernel on $X_{\infty}$ (Theorem 6.5). This heat kernel defines a semigroup (Theorem 6.6), corresponding to a Dirichlet form that satisfies the desired properties. The verification of Mosco convergence relies on the strong convergence of semigroups, as stated in the Mosco-Kuwae-Shioya’s theorem (Theorem 7.3).
In [KS03, Section 5], the authors show that, under the uniform Poincaré inequality and other geometric conditions, the Dirichlet forms converge in the Mosco topology. However, it remains unknown whether the Poincaré inequality $\mathrm{(PI)}$ is stable—i.e., whether it holds for the limit Dirichlet form. In [HKST15, Section 11.6], the stability of the $p$-Poincaré inequality for upper gradients is proved. Our results indicate that the uniform heat kernel bounds on varying metric measure Dirichlet spaces imply a heat kernel bound on the limit metric measure Dirichlet space. Building on the well-known results on heat kernel estimates [GHL15, Theorem 1.2], Theorem 1.2 in this paper establishes that the package of conditions $\mathrm{(V)}+\mathrm{(PI)}+\mathrm{(CSA)}$ remain stable under Gromov-Hausdorff convergence, where $\mathrm{(V)}$ represents the homogeneous volume growth estimate and $\mathrm{(CSA)}$ represents the cutoff Sobolev inequality (see [AB15] or [GHL15, p.1492] for definition).
This paper is organized as follows. In Section 2, we review the definition of pointed Gromov-Hausdorff convergence. In Section 3, we construct a doubling measure on the limit space and verify Theorem 1.2-(1). In Section 4, we utilize the Hilbert-Schmidt expansion of the heat kernel and estimate the eigenvalues and eigenfunctions. In Section 5, we construct the heat kernel with Dirichlet boundary conditions in the balls of the limit space. In Section 6, we obtain a heat kernel in the limit space and prove Theorem 1.2-(2). In Section 7, the Mosco convergence of Dirichlet forms, stated in Theorem 1.2-(3), is established. In Section 8, we apply Theorem 1.2 to reconstruct the diffusions on the Sierpiński carpet and Sierpiński gasket. We also provide an alternative proof in Theorem 8.4 that for any $\alpha\geq 1$ and $12\leq\beta\leq\alpha+$ 1, there exists a metric measure Dirichlet space that satisfies volume growth $r^{\alpha}$ and $\operatorname{HKE}(r^{\beta})$.
Notation.
In this paper, we use the following notation and conventions.
  1. (i)
    For $1\leq j\leq\infty$, $(x,r)\in X_{j}\times(0,\infty)$, we write the open ball in $X_{j}$ centered at $x$ with radii $r$
    $B_{j}(x,r):=\{z\in X_{j}:d_{j}(x,z)<r\}.$
  2. (ii)
    Let $X$ be a non-empty set. We define $\mathds{1}_{A}=\mathds{1}_{A}^{X}\in\mathbb{R}^{X}$ for $A\subset X$ by
    $\mathds{1}_{A}(x):=\mathds{1}_{A}^{X}(x):=\begin{cases}1&\mbox{if $x\in A$,}\\ 0&\mbox{if $x\notin A$.}\end{cases}$
  3. (iii)
    Let $X$ be a topological space. We set $C(X):=\{f\mid\textrm{$f:X\to\mathbb{R}$, $f$ is continuous}\}$, $C_{b}(X):=\{f\in C(X)\mid\textrm{$f$ is bounded}\}$, $C_{0}(X):=\{f\in C(X)\mid\textrm{$f$ vanishes at infinity}\}$ and
    $C_{c}(X):=\{f\in C(X)\mid\textrm{$X\setminus f^{-1}(0)$ has compact closure in$X$}\}.$
  4. (iv)
    For a measure $\mu$ on $(X,\mathcal{B}(X))$ and a mesureable function $f:X\rightarrow Y$, we write $f_{\#}\mu$ the push-forward measure defined by $f_{\#}\mu(A)=\mu(f^{-1}(A))$ for all measureable subset of $Y$.
  5. (v)
    $\mathbb{Q}_{+}:=\mathbb{Q}\cap(0,\infty)$ is the set of all positive rational numbers. :=$\mathbb{N}:=\{1,2,\ldots\}$, that is $0\notin\mathbb{N}$.
  6. (vi)
    We use the letters $C$, $c$ etc. to denote positive constant whose value is inessential and may change at each occurrence.
  7. (vii)
    For real valued quantities $f$ and $g$, if there exists an implicit constant $C\geq 1$ that depends on inessential parameters such that $f\leq Cg$ then we write $f\lesssim g$. We write $f\asymp g$ whenever $f\lesssim g$ and $g\lesssim f$.
2 Gromov-Hausdorff convergence
In this section, we review the basic theory of Gromov-Hausdorff distance and Gromov-Hausdorff convergence. The main reference is [HKST15, Chapter 11].
Let $(Z,d_{Z})$ be a metric space. We say $(Z,d_{Z})$ is proper if the closed balls $\{y\in Z:d_{Z}(z,y)\leq r\}$ is compact for all $(z,r)\in Z\times(0,\infty)$. We say $(Z,d_{Z})$ is a length space, if the distance between any two points equals the smallest possible length of curves connecting them. By [HKST15, Lemma 8.3.11], every proper length spaces are geodesic, i.e, any two points can be connected by a curve whose length matches the exact distance between them.
For any $\epsilon>0$ and any nonempty subset $A\subset X$, the $\epsilon$-neighborhood of $A$ is defined as the set
$N_{\epsilon}(A):=\left\{z\in Z:\mathrm{dist}_{Z}(z,A)<\epsilon\right\},$
where $\mathrm{dist}_{Z}$ is the distance under the metric $d_{Z}$. The Hausdorff distance between two nonempty subsets $A,B\subset Z$ is
$d_{\mathrm{H}}(A,B):=\inf\left\{\epsilon>0:A\subset N_{\epsilon}(B)\text{ and }B\subset N_{\epsilon}(A)\right\}.$
We define for each $x=\{x_{j}\}_{j=1}^{\infty},y=\{y_{j}\}_{j=1}^{\infty}\in\mathbb{R}^{\mathbb{N}}$ the quantities $\lVert x\rVert_{\infty}:=\sup_{1\leq j<\infty}|x_{j}|$ and
$d_{l^{\infty}}(x,y):=\lVert x-y\rVert_{\infty}=\sup_{1\leq j<\infty}|x_{j}-y_{j}|.$
Let
$l^{\infty}:=\left\{x=\{x_{j}\}_{j=1}^{\infty}\in\mathbb{R}^{\mathbb{N}}:\lVert x\rVert_{\infty}<\infty\right\},$
so that $(l^{\infty},d_{l^{\infty}})$ is a complete metric space. Recall that proper metric spaces are always separable. By the Fréchet embedding theorem [HKST15, p.101], every separable metric space $(X,d_{X})$ allows an isometric embedding $i:(X,d_{X})\to(l^{\infty},d_{l^{\infty}})$, i.e., $d_{l^{\infty}}(i(x_{1}),i(x_{2}))=d_{X}(x_{1},x_{2})$ for all $x_{1},x_{2}\in X$. The Gromov-Hausdorff distance between two separable metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ is defined as
$d_{\mathrm{GH}}(X,Y):=\inf d^{\infty}_{\mathrm{H}}(i(X),j(Y)),$
where $d^{\infty}_{\mathrm{H}}$ is the Hausdorff distance in $(l^{\infty},d_{l^{\infty}})$, and the infimum is taken over all isometric embeddings $i:(X,d_{X})\to(l^{\infty},d_{l^{\infty}})$ and $j:(Y,d_{Y})\to(l^{\infty},d_{l^{\infty}})$. For a sequence of separable metric spaces $\{(X_{n},d_{n}):1\leq n\leq\infty\}$, we write $(X_{n},d_{n})\xrightarrow{\mathrm{GH}}(X_{\infty},d_{\infty})$ if $d_{\mathrm{GH}}(X_{n},X_{\infty})\to 0$ as $n\to\infty$, and we say that the sequence $\{(X_{n},d_{n}):1\leq n<\infty\}$ Gromov–Hausdorff converges to $(X_{\infty},d_{\infty})$. It is known that the Gromov-Hausdorff distance $d_{\mathrm{GH}}$ gives a complete, separable, and contractible metric on the collection of all isometry classes of compact metric spaces [HKST15, Theorem 11.1.15].
As pointed out in [HKST15, Section 11.3], the Gromov-Hausdorff convergence is quite restrictive for the class of non-compact metric spaces. Therefore, the following concept of pointed Gromov–Hausdorff convergence is developed to deal with (possibly) non-compact spaces.
Definition 2.1 (Pointed Gromov-Hausdorff convergence).
A sequence of pointed separable metric spaces $\{(X_{j},d_{j},p_{j}):1\leq j<\infty\}$ is said to pointed Gromov–Hausdorff converge to a pointed separable metric space $(X_{\infty},d_{\infty},p_{\infty})$ if, for each $r>0$ and $0<\epsilon<r$, there exists an $i_{0}$ such that for each $i\geq i_{0}$ there is a map $f_{i}^{\epsilon}:B_{i}(p_{i},r)\rightarrow X_{\infty}$ satisfying:
  1. (1)
    $f_{i}^{\epsilon}(p_{i})=p_{\infty}$;
  2. (2)
    $|d(f_{i}^{\epsilon}(x),f_{i}^{\epsilon}(y)-d_{i}(x,y)|<\epsilon$ for all $x,y\in B_{i}(p_{i},r)$;
  3. (3)
    $B_{\infty}(p_{\infty},r-\epsilon)\subset N_{\epsilon}(f_{i}^{\epsilon}(B_{i}(p_{i},r)))$, where $N_{\epsilon}(A)$ is the $\epsilon$-neighborhood of $A$.
In this case, we write $(X_{n},d_{n},p_{n})\xrightarrow{\mathrm{p-GH}}(X_{\infty},d_{\infty},p_{\infty})$.
By [HKST15, Proposition 11.3.5], for bounded separable spaces, the notions of Gromov-Hausdorff convergence and pointed Gromov-Hausdorff convergence coincide. Gromov’s compactness theorem in this context can be stated as follows: for every eventually proper sequence in a pointed totally bounded family of pointed length metric spaces, there exists a subsequence that pointed Gromov–Hausdorff converges to a proper length pointed metric space [HKST15, Proposition 11.3.12, Theorem 11.3.16].
In general, the Gromov-Hausdorff approximations $f_{i}^{\epsilon}$ introduced in Definition 2.1 may not be measurable with respect to the Borel $\sigma$-algebra. This lack of measurability brings difficulties when analyzing such approximations. However, as we will see later, under the volume doubling property, we can carefully construct $f_{i}^{\epsilon}$ to be measurable functions. To begin, we recall the following covering theorem, which is due to David [Dav88] and Chrits [Chr90].
Theorem 2.2 ([Chr90, Theorem 11]).
Let $(Z,d,\mu)$ be a separable, volume doubling metric measure space, that is, there exists $C_{Z}>0$ such that $\mu(B(z,2r))\leq C_{Z}\mu(B(z,r))$ for all $z\in Z$ and all $r>0$. Then there exists some index set $\{I_{k}\}_{k\in\mathbb{Z}}$, a countable collection of open subsets $\left\{Q_{k,j}\subset X\;\big{|}\;j\in I_{k},\ k\in\mathbb{Z}\right\}$, a countable set of points $\left\{z_{k,j}\in X\;\big{|}\;j\in I_{k},\ k\in\mathbb{Z}\right\}$, and constants $\delta\in(0,1)$, $0<a_{0}\leq a_{1}<\infty$, such that
  1. (1)
    For all $k\in\mathbb{Z}$,
    $\mu\left(Z{\bigg{\backslash}}\bigcup_{j\in I_{k}}Q_{k,j}\right)=0.$ (2.1)
  2. (2)
    $Q_{k,i}\cap Q_{k,j}=\emptyset$ for each $k\in\mathbb{Z}$ and $i\neq j\in I_{k}$.
  3. (3)
    For each $k\in\mathbb{Z}$ and $j\in I_{k}$, $B(z_{k,j},a_{0}\delta^{k})\subset Q_{k,j}\subset B(z_{k,j},a_{1}\delta^{k})$.
Moreover, the constants $\delta\in(0,1)$ and $0<a_{0}\leq a_{1}<\infty$ depend only on $C_{Z}$.
Proof.
The volume doubling property is equivalent to the definition of being homogeneous type in [Chr90, Definition 1]. By [Chr90, Theorem 11 (3.1), (3.4) and (3.5)], we can find cubes that satisfy (1) and (3). (2) follows from [Chr90, Lemma 15]. ∎
The following proposition is the main result of this section.
Proposition 2.3.
Assume (A4), (A5) and (A6). Then there exist two sequences of positive numbers $\{\epsilon_{n}\}_{n=1}^{\infty}$ and $\{R_{n}\}_{n=1}^{\infty}$ with $\epsilon_{n}\downarrow 0$ and $R_{n}\uparrow\infty$ as $n\rightarrow\infty$, and two sequences of maps $f_{n}:\overline{B_{{n}}(p_{{n}},R_{n})}\rightarrow{B_{\infty}(p_{\infty},R_{n})}$ and $g_{n}:\overline{B_{\infty}(p_{\infty},R_{n})}\rightarrow{B_{{n}}(p_{{n}},R_{n})}$, $n\geq 1$ such that the following are true:
  1. (1)
    For all $n\geq 1$, the maps $f_{n}$ and $g_{n}$ are Borel measurable with respect to the completion of Borel $\sigma$-algebra on $X_{n}$ and $X_{\infty}$, respectively.
  2. (2)
    For all $n\geq 1$,
    $f_{n}(\overline{B_{\infty}(p_{\infty},R_{n})})$ is a finite $(2\epsilon_{n})$-net in $B_{\infty}(p_{\infty},R_{n})$. (2.2)
  3. (3)
    For all $x,y\in\overline{B_{n}(p_{n},R_{n})}$,
    $\left|d_{\infty}(f_{n}(x),f_{n}(y))-d_{n}(x,y)\right|\leq\epsilon_{n},$ (2.3)
    $d_{n}(x,g_{n}(f_{n}(x)))\leq\epsilon_{n}.$ (2.4)
  4. (4)
    For all $x,y\in\overline{B_{\infty}(p_{\infty},R_{n})}$,
    $\left|d_{\infty}(x,y)-d_{n}(g_{n}(x),g_{n}(y))\right|\leq\epsilon_{n},$ (2.5)
    $d_{\infty}(x,f_{n}(g_{n}(x)))\leq\epsilon_{n}.$ (2.6)
  5. (5)
    For all $n\geq 1$,
    $f_{n}(p_{n})$ $=p_{\infty},$ (2.7)
    $g_{n}(p_{\infty})$ $=p_{n}.$ (2.8)
We call $f_{n}$ and $g_{n}$ as a family of $\epsilon_{n}-$approximate isometries.
Proof.
Let $\{\epsilon_{n}\}_{n=1}^{\infty}$ and $\{R_{n}\}_{n=1}^{\infty}$ be two positive sequences that will be chosen later, with $\epsilon_{n}\downarrow 0$ and $R_{n}\uparrow\infty$ as $n\rightarrow\infty$.
Recall that a closed ball in a length space is the closure of the respective open ball. By the proof of [HKST15, Proposition 11.3.12], for each $R>1$, we have
$\overline{B_{n}(p_{n},R)}\xrightarrow{\mathrm{GH}}\overline{B_{\infty}(p_{\infty},R)}\text{ as $n\rightarrow\infty$},$
i.e.,
$d_{\text{GH}}(\overline{B_{n}(p_{n},R)},\overline{B_{\infty}(p_{\infty},R)})\rightarrow 0\text{ as $n\rightarrow\infty$.}$
By [Jan17, Lemma 2.8, Proposition 2.9], there are monotone positive sequences $R_{n}\uparrow\infty$ and $s_{n}\downarrow 0$ such that $d_{\text{GH}}(B_{{n}}(p_{{n}},R_{n}),B_{\infty}(p_{\infty},R_{n}))\leq s_{n}\rightarrow 0$ as $n\rightarrow\infty$. For each $n\geq 1$, we apply [HKST15, Proposition 11.1.9] and assume that $(B_{n}(p_{n},R),d_{n})$ and $(B_{\infty}(p_{\infty},R),d_{\infty})$ are isometric embedded into a common metric space $(Z,d_{Z})$. We do not distinguish $(B_{n}(p_{n},R),d_{n})$ and $(B_{\infty}(p_{\infty},R),d_{\infty})$ with their isometric images, respectively. From [HKST15, Proposition 11.1.11 and (11.1.12)] we have
$d^{Z}_{\text{H}}(\overline{B_{n}(p_{n},R_{n})},\overline{B_{\infty}(p_{\infty},R_{n})})+d_{Z}(p_{n},p_{\infty})\leq 5s_{n}.$ (2.9)
We first define the following:
  1. (i)
    For each $n\geq 1$ we decompose the volume doubling space $(X_{n},d_{n},m_{n})$ according to Theorem 2.2 and obtain a index set $\{I^{(n)}_{k}\}_{k\in\mathbb{Z}}$, a family of open sets $\left\{Q_{k,j}^{(n)}\subset X_{n}\,\big{|}\,j\in I^{(n)}_{k},k\in\mathbb{Z}\right\}$, a set of points $\left\{z_{k,j}^{(n)}\in X_{n}\,\big{|}\,j\in I^{(n)}_{k},k\in\mathbb{Z}\right\}$ and constants $\delta\in(0,1)$ and $0<a_{0}\leq a_{1}$ that are independent of $n$. Define
    $J_{k}^{(n)}:=\{j\in I^{(n)}_{k}:Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\neq\emptyset\}.$
    For each $j\in J_{k}^{(n)}$, if $z_{k,j}^{(n)}\in Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}$, we define $\widetilde{z}_{k,j}^{(n)}$ to be $z_{k,j}^{(n)}$, and if otherwise we define $\widetilde{z}_{k,j}^{(n)}$ to be any arbitrary (but fixed) point in $Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}$.
  2. (ii)
    Since $\overline{B_{\infty}(p_{\infty},R)}$ is compact, thus totally bounded, we can find a finite $s_{n}$-net of $\overline{B_{\infty}(p_{\infty},R)}$ with cardinality $N_{n}$, say $\{w_{j}\}_{j=1}^{N_{n}}$. We slightly modify this net so that $\{w_{j}\}_{j=1}^{N_{\epsilon}}\subset{B_{\infty}(p_{\infty},R)}$. For each $z\in\overline{B_{n}(p_{n},R)}$, by (2.9) we are able to find $w_{z}\in\{w_{j}\}_{j=1}^{N_{n}}$ such that $d_{Z}(z,w_{z})\leq 6s_{n}$.
We now define $f_{k,n}:\overline{B_{n}(p_{n},R)}\to\{w_{j}\}_{j=1}^{N_{n}}\cup\{p_{\infty}\}\subset B_{\infty}(p_{\infty},R_{n})$ by
$f_{k,n}(z):=\begin{cases}p_{\infty}\;&\text{ if }z=p_{n}\\ w_{\widetilde{z}_{k,j}^{(n)}}\;&\text{ if }z\in\left(Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\right)\setminus\{p_{n}\},\ j\in J_{k}^{(n)}\\ w_{z}\quad&\text{ if }z\in\left(\overline{B_{n}(p_{n},R)}\setminus\bigcup_{j\in J_{k}^{(n)}}Q_{k,j}^{(n)}\right)\setminus\{p_{n}\}.\end{cases}$ (2.10)
It is easy to verify, using (2.9), the definition of $z\mapsto w_{z}$ in (ii) and triangle inequality that
$d_{Z}(f_{n,k}(z),z)\leq d_{Z}(f_{n,k}(z),w_{z})+d_{Z}(w_{z},z)\leq 2a_{1}\delta^{k}+6s_{n}\text{ for all $z\in\overline{B_{n}(p_{n},R)}$}.$ (2.11)
Therefore for $z_{1},z_{2}\in\overline{B_{n}(p_{n},R)}$,
|$\phantom{\leq\ }\absolutevalue{d_{\infty}(f_{n,k}(z_{1}),f_{n,k}(z_{2}))-d_{n}(z_{1},z_{2})}$ | (2.12)
=|$=\absolutevalue{d_{Z}(f_{n,k}(z_{1}),f_{n,k}(z_{2}))-d_{Z}(z_{1},z_{2})}$ | (2.13)
$\leq d_{Z}(f_{n,k}(z_{1}),z_{1})+d_{Z}(f_{n,k}(z_{2}),z_{2})\leq 12(a_{1}\delta^{k}+s_{n}).$ (2.14)
For each $w\in\overline{B_{\infty}(p_{\infty},R_{n})}$, we can find a $v_{w}\in\overline{B_{n}(p_{n},R)}$ such that $d_{Z}(w,v_{w})\leq 6s_{n}$. Since the ball $B_{n}(v_{w},s_{n})$ must intersect one of the set $\{Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\}_{j\in J_{k}^{(n)}}$ as the measure $m_{n}$ has full support, we can choose $\widetilde{v}_{z}\in\bigcup_{j\in J_{k}^{(n)}}\left(Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\right)\setminus\{p_{n}\}$ so that $d_{Z}(w,\widetilde{v}_{w})\leq 7s_{n}$. Therefore
$d_{\infty}(f_{k,n}(\widetilde{v}_{w}),w)$ $=d_{Z}(f_{k,n}(\widetilde{v}_{w}),w)$ (2.15)
$\leq d_{Z}(f_{k,n}(\widetilde{v}_{w}),\widetilde{v}_{w})+d_{Z}(\widetilde{v}_{w},w)$ (2.16)
(2.11)${\leq}12(a_{1}\delta^{k}+s_{n})+s_{n}\leq 24(a_{1}\delta^{k}+s_{n})$ ) (2.17)
so
$f_{k,n}(\overline{B_{\infty}(p_{\infty},R_{n})})$ is a finite $24(a_{1}\delta^{k}+s_{n})$-net in $B_{\infty}(p_{\infty},R_{n})$. (2.18)
Now let $f_{n}:=f_{n,n}$ and $\epsilon_{n}=12(a_{1}\delta^{n}+s_{n})$. Since every subset of $\left(\overline{B_{n}(p_{n},R)}\setminus\bigcup_{j\in J_{k}^{(n)}}Q_{k,j}^{(n)}\right)$ is in the completion of Borel $\sigma$-algebra on $\overline{B_{n}(p_{n},R)}$, we see that (1) holds for $f_{n}$. (2.2), (2.3) and (2.7) follows from (2.18), (2.14) and the definition of $f_{n}$, respectively.
From Proposition 3.2, whose proof depends only on existence of $f_{n}$, (2.3), (2.2) and (2.7), we see that $(X_{\infty},d_{\infty})$ admits a volume doubling measure. Thus we can apply Theorem 2.2 again to limit space $(X_{\infty},d_{\infty})$ and construct $g_{n}$, in the same way as we construct $f_{n}$, so that (1), (2.5) and (2.8) holds. By (2.3), (2.5) and triangle inequality,
$d_{n}(x,g_{n}(f_{n}(x)))\leq d_{Z}(x,f_{n}(x))+d_{Z}(f_{n}(x),g_{n}(f_{n}(x)))\leq\epsilon_{n},$
which gives (2.4). (2.6) holds by the same argument. ∎
The following lemma, concerning the convergence of diameters, will be used in subsequent estimates. Its proof is a direct adaptation of [Jan17, Proposition 2.13].
Lemma 2.4.
Under (A6), we have
$\mathrm{diam}(X_{j},d_{j})\to\mathrm{diam}(X_{\infty},d_{\infty})\text{ as $j\to\infty$}.$
3 Doubling measure on limit space
In this section, we assume $(X_{j},d_{j},p_{j})\to(X_{\infty},d_{\infty},p_{\infty})$ under the pointed Gromov-Hausdorff convergence. Let $\{f_{n}\}$ and $\{g_{n}\}$ be the family of $\epsilon_{n}$-approximate isometries in Proposition 2.3. The goal in this section is to construct a Radon measure $m_{\infty}$ on $X_{\infty}$ that satisfies (1.9).
Lemma 3.1 ([HKST15, Proposition 11.5.3]).
Let $(Z,d)$ be a length space and $\mu$ be a volume doubling measure on $X$ with constant $C_{\mathrm{v}}$. Then for every $z\in Z$ and $0\leq r<R$,
$\mu\left(B(z,R)\setminus\overline{B(z,r)}\right)\leq C\left(\frac{R-r}{R}\right)^{\gamma}\mu(B(z,R)),$
where
$\gamma=\frac{\log\left({1+C_{\mathrm{v}}^{-4}}\right)}{\log 3}\text{ and }C=6^{\gamma}.$
The following proposition includes the proof of Theorem 1.2-(1).
Proposition 3.2.
Assume (A4), (A5), and (A6). There exists a Radon measure $m_{\infty}$ on $(X_{\infty},d_{\infty})$ such that
  1. (1)
    the volume growth satisfies (1.9).
  2. (2)
    along some subsequence of $\{(X_{n},d_{n},p_{n})\}_{n=1}^{\infty}$,
    $\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\right)\Big{|}_{B_{\infty}(p_{\infty},R)}\rightarrow m_{\infty}|_{B_{\infty}(p_{\infty},R)}\quad\text{weakly as }n\rightarrow\infty,$
    for every $R\in\mathbb{Q}_{+}\cup\{R_{n}\}_{n=1}^{\infty}$, in the sense that, for all $\phi\in C(B_{\infty}(p_{\infty},R))$ that is bounded and uniformly continuous,
    $\int_{B_{\infty}(p_{\infty},R)}\phi\mathop{}\!\mathrm{d}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\rightarrow\int_{B_{\infty}(p_{\infty},R)}\phi\mathop{}\!\mathrm{d}m_{\infty}\text{ as }n\rightarrow\infty.$ (3.1)
Proof.
The proof consists of four steps. For each $R>0$, we will construct a measure on the ball $B_{\infty}(p_{\infty},R)$, and then unify these measures to get the desired globally defined measure.
  1. Step 1.
    For the first step, we fix the $0<R<\infty$. Let $\{\epsilon_{n}\}$ and $\{R_{n}\}$ be in Proposition 2.3. Since $R_{n}\rightarrow\infty$, we assume that $R_{n}>R+1$. By the triangle inequality, (2.3) and (2.7),
    $B_{n}(p_{n},R-\epsilon_{n})\subset\left\{x\in\overline{B_{n}(p_{n},R_{n})}\;\Big{|}\;f_{n}(x)\in B_{\infty}(p,R)\right\}\subset B_{n}(p_{n},R+\epsilon_{n}).$
    Thus
[…]
    We may choose a subsequence such that along this subsequence, the limit
    ${\lim_{n\rightarrow\infty}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)(B_{\infty}(p,R))}$
    exists and is positive.
  2. Step 2.
    For this $R$ and along this subsequence, we claim that $\left\{\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\right)\Big{|}_{B_{\infty}(p_{\infty},R)}\right\}$, as a set of measures on $B_{\infty}(p_{\infty},R)$, is tight, that is, for every $\epsilon>0$, there exists a compact subset $K\subset B_{\infty}(p_{\infty},R)$ such that
    $\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\left(B_{\infty}(p_{\infty},R)\setminus K\right)\leq\epsilon\quad\text{for all}n\geq 1.$
    In fact, for any given $\epsilon>0$, we can choose $r_{\epsilon}<R$ and $N_{\epsilon}\in\mathbb{N}_{+}$ such that, for all $r\in(r_{\epsilon},R)$ and $n\geq N_{\epsilon}$, by triangle inequality,
    $\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\left(B_{\infty}(p_{\infty},R)\setminus\overline{B_{\infty}(p_{\infty},r)}\right)$
    $=$ $m_{n}\left(\left\{x\in\overline{B_{n}(p_{n},R_{n})}\;\Big{|}\;f_{n}(x)\in B_{\infty}(p_{\infty},R)\setminus\overline{B_{\infty}(p_{\infty},r)}\right\}\right)$
    $\leq$ $m_{n}\left(B_{n}(p_{n},R+2\epsilon_{n})\setminus\overline{B_{n}(p_{n},r-2\epsilon_{n})}\right)$
    $\leq$ $6^{\gamma}C_{\mathrm{v}}(R-r+4\epsilon_{n})^{\gamma}(R+2\epsilon_{n})^{-\gamma}V_{u}(R+2\epsilon_{n})$
    $\leq$ $6^{\gamma}C_{\mathrm{v}}(R-r_{\epsilon}+4\epsilon_{n})^{\gamma}(R+2\epsilon_{n})^{-\gamma}V_{u}(R+2\epsilon_{n})<\epsilon$
    Since $\left\{\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\right)\Big{|}_{B_{\infty}(p_{\infty},R)}\right\}_{n=1}^{N_{\epsilon}}$ is a finite set and by (2.2) the elements in the image of $f_{n}$ is also finite for each $n$, we can choose a $r^{\prime}_{\epsilon}<R$ such that, for all $r\in(r^{\prime}_{\epsilon},R)$, $f_{n}^{-1}\left(B_{\infty}(p_{\infty},R)\setminus\overline{B_{\infty}(p_{\infty},r)}\right)=\emptyset$, $1\leq n\leq N_{\epsilon}$. Hence, for such $r$,
    $\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\left(B_{\infty}(p_{\infty},R)\setminus\overline{B_{\infty}(p_{\infty},r)}\right)=0,\quad 1\leq n\leq N_{\epsilon}$
    Thus for any $r\in(\max(r^{\prime}_{\epsilon},r_{\epsilon}),R)$,
    $\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\left(B_{\infty}(p_{\infty},R)\setminus\overline{B_{\infty}(p_{\infty},r)}\right)<\epsilon,\text{ for all }n\geq 1.$
    which means that $\left\{\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\right)\Big{|}_{B_{\infty}(p_{\infty},R)}\right\}_{n=1}^{\infty}$ is tight. We then apply the non-compact version of Prokhorov’s theorem [Tao10, Exercise 1.10.29] and a well-known fact that, if the vague convergent limit is a probability measure then the convergence holds in the weak sense [Bil99, Problem 5.10], there exist a non-zero probability measure $\widetilde{m_{R}}$ on $B_{\infty}(p_{\infty},R)$ such that, for some subsequence,
    $\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)(B_{\infty}(p,R))\right)^{-1}\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\right)\Big{|}_{B_{\infty}(p_{\infty},R)}\rightarrow\widetilde{m_{R}}\text{ weakly as }n\rightarrow\infty.$
    We now write $m_{R}=\lim_{n\rightarrow\infty}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)(B_{\infty}(p,R))\cdot\widetilde{m_{R}}.$ By a standard diagonal argument, we can extract a common subsequence of $\mathbb{N}$ such that
    $\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\right)\Big{|}_{B_{\infty}(p_{\infty},R)}\rightarrow m_{R}\quad\text{weakly, for all }R\in\mathbb{Q}_{+}\cup\{R_{n}\}_{n=1}^{\infty}.$ (3.2)
  3. Step 3.
    If $R_{1}\neq R_{2}$, let us show that $m_{R_{1}}$ and $m_{R_{2}}$ agree on their common domain. Without loss of generality, assume $R_{1}<R_{2}$. It is straightforward to see that
    $\left(\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\right)\Big{|}_{B_{\infty}(p_{\infty},R_{1})}\rightarrow m_{R_{2}}|_{B_{\infty}(p_{\infty},R_{1})}\quad\text{weakly},$
    by the uniqueness of weak convergence, we conclude that $m_{R_{2}}|_{B_{\infty}(p_{\infty},R_{1})}=m_{R_{1}}.$ We then can define a set function $m_{\infty}$ on $(X_{\infty},d_{\infty})$ by
    $m_{\infty}(E)=\lim_{\mathbb{Q}_{+}\ni R\rightarrow\infty}m_{R}\left(E\cap B_{\infty}(p_{\infty},R)\right),\quad\text{for all Borel subset of }(X_{\infty},d_{\infty}).$
    The limit exists, as the sequence is monotone increasing. It is direct to verify by [HKST15, Proposition 3.3.44] that $m_{\infty}$ is a Radon measure.
  4. Step 4.
    The final step is to verify that $m_{\infty}$ satisfies (1.9). In fact, for any $x\in X_{\infty}$ and $0<r<\mathrm{diam}(X_{\infty},d_{\infty})$, we may choose $n$ large enough such that $B_{\infty}(x,r)\subset B_{\infty}(p_{\infty},R_{n})$ and $r+{4}\epsilon_{n}<\mathrm{diam}(X_{n},d_{n})$ by Lemma 2.4. Since every metric space is normal, by the Urysohn’s lemma, there exist functions $\phi_{1}\in C_{c}(B_{\infty}(x,r);[0,1])$ with $\phi_{1}=1$ on $B_{\infty}(x,r-\epsilon_{n})$, and $\phi_{2}\in C_{c}(B_{\infty}(x,r+\epsilon_{n});[0,1])$ with $\phi_{2}=1$ on $B_{\infty}(x,r)$. By (2.2), we can find a $x_{n}\in B_{n}(p_{n},R_{n})$ such that $d_{\infty}(f_{n}(x_{n}),x)\leq 2\epsilon_{n}$. According to triangle inequality
    $C_{\mathrm{v}}^{-1}V_{n}(r-4\epsilon_{n})$ $\leq m_{n}\left(B_{n}\left(x_{n},r-4\epsilon_{n}\right)\right)$
    $\leq\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\left(B_{\infty}\left(x,r-\epsilon_{n}\right)\right)$
    $\leq\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)(B_{\infty}(x,r+\epsilon_{n}))$
    $\leq m_{n}(B_{n}(x_{n},r+4\epsilon_{n}))\leq C_{\mathrm{v}}V_{n}(r+4\epsilon_{n}).$
    Hence by the convergence of $V_{n}$ we have
    $C_{\mathrm{v}}^{-1}V_{\infty}(r)$ $\leq\liminf_{n\rightarrow\infty}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)(B_{\infty}(x,r-\epsilon_{n}))$
    $\leq\lim_{n\rightarrow\infty}\int_{B_{\infty}(x,r)}\phi_{1}\mathop{}\!\mathrm{d}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)$
    $\leq m_{R}(B_{\infty}(x,r))=m_{\infty}(B_{\infty}(x,r))$
    $\leq\lim_{n\rightarrow\infty}\int_{B_{\infty}(x,r)}\phi_{2}\mathop{}\!\mathrm{d}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)\leq C_{\mathrm{v}}V_{\infty}(r),$
    which gives (1.9). The weak convergence of measures is given by (3.2).
The following concept of convergence of the $L^{2}$ space is a special case of convergence of the Hilbert space considered in [KS03, Section 2.2].
Definition 3.3.
Let $\left\{(X_{n},m_{n}):1\leq n\leq\infty\right\}$ be a sequence of measurable space. We say that the Hilbert spaces $L^{2}(X_{n},m_{n})$ converge to $L^{2}(X_{\infty},m_{\infty})$ if there exists a dense subspace $\mathcal{C}\subset L^{2}(X_{\infty},m_{\infty})$ and a sequence of linear operators $\upphi_{n}:\mathcal{C}\to L^{2}(X_{n},m_{n})$ with the following property:
$\lim_{n\to\infty}\lVert\upphi_{n}u\rVert_{L^{2}(X_{n},m_{n})}=\lVert u\rVert_{L^{2}(X_{\infty},m_{\infty})},\ \forall u\in\mathcal{C}.$ (3.3)
Theorem 3.4.
Assume (A4), (A5), (A6) and let $m_{\infty}$ be given by Proposition 3.2. Then $L^{2}(X_{n},m_{n})$ converges to $L^{2}(X_{\infty},m_{\infty})$.
Proof.
Let $\mathcal{C}:=C_{c}(X_{\infty})$. Define $\upphi_{n}:\mathcal{C}\to L^{2}(X_{n},m_{n})$ by
$\upphi_{n}u(x)=\begin{dcases}(u\circ f_{n})(x)\ \quad&x\in\overline{B_{n}(p_{n},R_{n})}\\ 0\ \quad&x\in X_{n}\setminus\overline{B_{n}(p_{n},R_{n})},\end{dcases}$ (3.4)
For $u\in\mathcal{C}$, we assume that $\mathrm{supp}(u)\in B_{\infty}(p_{\infty},R)$ for some $R\in\mathbb{Q}_{+}$. Then by (3.1)
$\lVert\upphi_{n}u\rVert_{L^{2}(X_{n},m_{n})}^{2}$ $=\int_{X_{n}}(\upphi_{n}u)^{2}\mathop{}\!\mathrm{d}m_{n}=\int_{\overline{B_{n}(p_{n},R_{n})}}(u\circ f_{n})(x)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$ (3.5)
$=\int_{B_{\infty}(p_{\infty},R)}u^{2}\mathop{}\!\mathrm{d}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R)}}\right)$ (3.6)
$\to\int_{B_{\infty}(p_{\infty},R)}u^{2}\mathop{}\!\mathrm{d}m_{\infty}=\lVert u\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}$ (3.7)
as $n\to\infty$. ∎
4 Estimates on heat kernel and eigenfunctions
In this section, we will estimate the growth of eigenvalues and the eigenfunctions of Dirichlet forms on balls with Dirichlet boundary conditions. Throughout this section, we assume that all the conditions in Theorem 1.2 are satisfied.
We begin by briefly reviewing some standard notation related to Dirichlet forms and refer the reader to [FOT11] for a more detailed discussion. Given a metric measure space $(X,d,m)$, we say that $(\mathcal{E},\mathcal{F})$ is a regular, symmetric Dirichlet form on $L^{2}(X,m)$, if it satisfies the following conditions:
  1. (i)
    $\mathcal{F}$ is a dense linear subspace of $L^{2}(X,m)$;
  2. (ii)
    $\mathcal{E}$ is a non-negative definite symmetric bilinear form on $\mathcal{F}\times\mathcal{F}$;
  3. (iii)
    $\mathcal{F}$ is a Hilbert space with inner product $\mathcal{E}_{1}:=\mathcal{E}+\langle\cdot,\cdot\rangle_{L^{2}(X,m)}$;
  4. (iv)
    for every $u\in\mathcal{F}$, $v:=(0\vee u)\wedge 1\in\mathcal{F}$ and $\mathcal{E}(v,v)\leq\mathcal{E}(u,u)$;
  5. (v)
    $\mathcal{F}\cap C_{c}(X)$ is dense both in $(\mathcal{F},\sqrt{\mathcal{E}_{1}})$ and in $(C_{c}(X),\lVert\cdot\rVert_{\sup})$.
If, in addition, $(\mathcal{E},\mathcal{F})$ satisfies
  1. (vi)
    $\mathcal{E}(f,g)=0$ for any $f,g\in\mathcal{F}$ with $\mathrm{supp}_{m}(f)$, $\mathrm{supp}_{m}(g)$ compact, and $\mathrm{supp}_{m}(f-a\mathds{1}_{X})\cap\mathrm{supp}_{m}(g)=\emptyset$ for some $a\in\mathbb{R}$. Here, $\mathrm{supp}_{m}(f)$ denotes the support of the measure $|f|\mathop{}\!\mathrm{d}m$.
then we call $(\mathcal{E},\mathcal{F})$ a regular, strongly local and symmetric Dirichlet form. Given an open set $D\subset X$, we define another Dirichlet form $(\mathcal{E}^{D},\mathcal{F}^{D})$ by letting $\mathcal{F}^{D}$ be the closure of $\mathcal{F}\cap C_{c}(D)$ under the norm of $\mathcal{E}_{1}$, this ensures that $\mathcal{F}^{D}$ is a linear space of $\mathcal{F}$. For $u,v\in\mathcal{F}^{D}\subset\mathcal{F}$, we define $\mathcal{E}^{D}(u,v):=\mathcal{E}(u,v)$. It is known that $(\mathcal{E}^{D},\mathcal{F}^{D})$ is also a regular symmetric Dirichlet form on $L^{2}(D,{\left.\kern-1.2ptm\vphantom{\big{|}}\right|_{D}})$ [CF12, Theorem 3.3.9], and is called the part of $(\mathcal{E},\mathcal{F})$ on $D$.
Let $\{P_{t}:L^{2}(X,m)\to L^{2}(X,m)\}_{t>0}$ denote the semigroup corresponding to $(\mathcal{E},\mathcal{F})$. We say $(\mathcal{E},\mathcal{F})$ is conservative if $P_{t}\mathds{1}_{X}=\mathds{1}_{X}$ for every $t>0$. By [BGK12, Theorem 3.1], if $(\mathcal{E},\mathcal{F})$ satisfies the heat kernel estimates stated in Definition 1.1, and if $(X,d,m)$ satisfies the volume doubling property, then the heat kernels of $(\mathcal{E},\mathcal{F})$ and of $(\mathcal{E}^{D},\mathcal{F}^{D})$ are continuous.
Returning to our setting, consider a sequence of metric measure spaces $(X_{n},d_{n},m_{n})$ and, for each $n\geq 1$, a regular, strongly local symmetric Dirichlet form $(\mathcal{E}_{n},\mathcal{F}_{n})$ on $L^{2}(X_{n},m_{n})$. For each $R>0$, we write $(\mathcal{E}_{n}^{R},\mathcal{F}_{n}^{R}):=(\mathcal{E}_{n}^{B_{n}(p_{n},R)},\mathcal{F}_{n}^{B_{n}(p_{n},R)})$ as the part of $(\mathcal{E}_{n},\mathcal{F}_{n})$ on $B_{n}(p_{n},R)$ for brevity. The semigroup corresponding to $(\mathcal{E}_{n}^{R},\mathcal{F}_{n}^{R})$ is denoted by $\{P^{(n,R)}_{t}\}_{t>0}$, and its heat kernel by $p^{(n,R)}(t,x,y)$ for $t>0$ and $x,y\in B_{n}(p_{n},R)$. For convenience, we set $p^{(n,R)}(t,x,y)=0$ if $x$ or $y$ is outside $B_{n}(p_{n},R)$.
By [Dav89, Theorem 2.1.4], the generator of $(\mathcal{E}_{n}^{R},\mathcal{F}_{n}^{R})$ has a discrete spectrum $\{\lambda_{n,j}^{(R)}\}_{j=1}^{\infty}$ (written in increasing order and repeated according to multiplicity). Let $\{\phi^{(R)}_{n,j}\}_{j=1}^{\infty}$ be the corresponding eigenfunction normalized by $\lVert\phi_{n,j}\rVert_{L^{2}(B_{n}(p_{n},R),m_{n})}=1$. Moreover, the following Hilbert-Schmidt expansion holds:
$p^{(n,R)}(t,x,y)=\sum_{j=1}^{\infty}\exp(-\lambda_{n,j}^{(R)}t)\phi^{(R)}_{n,j}(x)\phi^{(R)}_{n,j}(y)\quad\text{for all }t>0\text{ and }x,y\in B_{n}(p_{n},R),$ (4.1)
where this series converges uniformly on $[\delta,\infty)\times B_{n}(p_{n},R)\times B_{n}(p_{n},R)$ for any $\delta>0$ and $n\geq 1$. We first estimate the growth of eigenvalues, which requires the following capacity condition:
Lemma 4.1.
There exist constants $\kappa_{\mathrm{cap}}\in(0,1)$ and $C_{\mathrm{cap}}>1$ such that, for any $n\geq 1$ and any ball $B_{n}=B_{n}(x,r)$ in $(X_{n},d_{n})$ of radius $r>0$, there exists a continuous cut-off function $\phi\in\mathcal{F}_{n}$ with $0\leq\phi\leq 1$, $\phi=1$ in an open neighborhood of $\kappa_{\mathrm{cap}}B_{n}$, $\mathrm{supp}(\phi)\subset B_{n}$, and
$\mathcal{E}_{n}(\phi,\phi)\leq C_{\mathrm{cap}}\frac{V_{n}(r)}{\Psi_{n}(r)}.$ (4.2)
Proof.
As the proof of [GHL15, Theorem 1.2] is quantitative with the constant, we see that the capacity condition $(\text{cap})_{\beta}$ in [GHL15] holds for all $(\mathcal{E}_{n},\mathcal{F}_{n})$ with the same constant $\kappa$ and $C$, which implies (4.2). ∎
Proposition 4.2.
There exists a constant $C_{1}\geq 1$, independent of $n$ and $R$, such that
[…]
Therefore
[…]
Proof.
We use the max-min principle [Sch12, Exercise 12.4.2]. Given an integer $j$ sufficiently large, we can find a disjoint family of $j$ balls $B_{n}(x_{i},r)\subset B_{n}(p,R)$, where $r=R/(100j)$. In fact, we can find two points with a distance greater than $R/2$ and choose one of the geodesics that connect those two points. We can pick points along this geodesic such that each two points have a distance greater than $R/(25j)$, then choose the balls centered at these points with radius $r=R/(100j)$. Applying Lemma 4.1 on these balls,
$\mathcal{E}_{n}^{R}(\phi_{i},\phi_{i})\lesssim\frac{V_{n}(r)}{\Psi_{n}(r)}.$
Thus by the variational characterization of the $j$-th eigenvalue $\lambda_{n,j}^{(R)}$, we see that such that for $\phi=\sum_{i=1}^{j}a_{i}\phi_{i}\in\mathrm{span}(\phi_{i},1\leq i\leq j)$ where $(a_{1},\ldots,a_{j})\in\mathbb{R}^{j}\setminus\{0\}$,
$\frac{\mathcal{E}^{R}_{n}(\phi,\phi)}{\lVert\phi\rVert_{L^{2}(B_{n}(p_{n},R))}^{2}}\lesssim\frac{\sum_{i=1}^{j}a_{i}^{2}\mathcal{E}^{R}(\phi_{i},\phi_{i})}{\sum_{i=1}^{j}a_{i}^{2}m(B_{n}(x_{i},\kappa_{\mathrm{cap}}r))}\lesssim\frac{V_{n}(r)/\Psi_{n}(r)}{V_{n}(\kappa_{\mathrm{cap}}r)}\lesssim\frac{1}{\Psi_{n}(R/j)}.$
Taking supremum over all $(a_{1},\ldots,a_{j})\in\mathbb{R}^{j}\setminus\{0\}$, we see by the max-min principle that
$\lambda_{n,j}^{(R)}=\inf_{H\subset\mathcal{F}_{n}:\dim H=j}\sup_{\phi\in H}\frac{\mathcal{E}^{R}_{n}(\phi,\phi)}{\lVert\phi\rVert_{L^{2}(B_{n}(p_{n},R))}^{2}}\leq\sup_{\phi\in\mathrm{span}(\phi_{i},1\leq i\leq j)}\frac{\mathcal{E}^{R}_{n}(\phi,\phi)}{\lVert\phi\rVert_{L^{2}(B_{n}(p_{n},R))}^{2}}\lesssim\frac{1}{\Psi_{n}(R/j)}.$
Remark 4.3.
If the metric space $(X_{n},d_{n})$ is bounded, that is, $\mathrm{diam}(X_{n},d_{n})<\infty$, and if $R>\mathrm{diam}(X_{n},d_{n})$, then $\lambda_{n,1}^{(R)}=0$, so the lower bound of $\lambda_{n,j}^{(R)}$ in (4.3) cannot be expected for $R$ very large. On the other hand, by the Faber-Krahn inequality (see [GHL15] for a definition), there exists $C_{2},C_{3}>1$ such that $\lambda_{n,1}^{(R)}\geq{C_{3}^{-1}(\Psi_{n}(R))^{-1}}$ for all $0<R<C^{-1}_{2}\mathrm{diam}(X_{n},d_{n})$.
Remark 4.4.
If $\Psi(r)=r^{\beta}$ for some $\beta>1$ and if the space is $\alpha$-regular, namely $V(r)\asymp r^{\alpha}$, then (4.3) gives $\lambda_{j}\lesssim j^{\beta}$, which is weaker than the classic Weyl’s law $\lambda_{j}\lesssim j^{\beta/\alpha}$ as by a recent result, $12\leq\beta\leq\alpha+$ 1 whenever the heat kernel estimate in Definition 1.1 holds [Mur24].
We also have the following $L^{\infty}$ estimates of eigenfunctions.
Proposition 4.5.
There exists a constant $C>0$, such that for each $n,j\geq 1$,
$\left\lVert\phi_{n,j}^{(R)}\right\rVert_{L^{\infty}\left(B_{n}(p_{n},R)\right)}\leq C\inf_{t\in(0,\infty)}\frac{e^{\lambda_{n,j}^{(R)}t}}{V_{n}(\Psi_{n}^{-1}(t))^{1/2}}.$ (4.4)
Proof.
Since $\phi_{n,j}^{(R)}$ is also the eigenvalue of the bounded operator $P^{(n,R)}_{t}$ with eigenvalue $\exp(-\lambda_{n,j}^{(R)}t)$, we have by the Cauchy-Schwarz’s inequality that
$\left|\phi_{n,j}^{(R)}(x)\right|$ $=\left|e^{\lambda_{n,j}^{(R)}t}P^{(n,R)}_{t}\phi_{n,j}^{(R)}(x)\right|=\left|e^{\lambda_{n,j}^{(R)}t}\int_{B_{n}(p_{n},R)}p^{(n,R)}(t,x,y)\phi_{n,j}^{(R)}(y)\mathop{}\!\mathrm{d}m_{n}(y)\right|$
$\leq{e^{\lambda_{n,j}^{(R)}t}}p^{(n,R)}(2t,x,x)^{1/2}\left\lVert\phi_{n,j}^{(R)}\right\rVert_{L^{2}\left(B_{n}(p_{n},R)\right)}$
$\lesssim\frac{e^{\lambda_{n,j}^{(R)}t}}{V_{n}(\Psi_{n}^{-1}(2t))^{1/2}}\lesssim\frac{e^{\lambda_{n,j}^{(R)}t}}{V_{n}(\Psi_{n}^{-1}(t))^{1/2}}.$
By taking the infimum over $t\in(0,\infty)$ we obtain (4.4). ∎
We need the Hölder estimates of the heat kernel and eigenfunctions to pass to the limit. To formulate it, we first define a metric on $X_{n}\times X_{n}$, with some abuse of notation, by
$d_{n}\left((x_{1},y_{1}),(x_{2},y_{2})\right):=\max(d_{n}(x_{1},x_{2}),d_{n}(y_{1},y_{2})).$
Under this metric, the ball in $X_{\infty}\times X_{\infty}$ centered at $(x,y)$ with radius $r$ equals $B_{\infty}(x,r)\times B_{\infty}(y,r)$, which is totally bounded.
Proposition 4.6.
For all $n\geq 1$, $j\geq 1$, $R>0$ and $t>0$, the Dirichlet heat kernels $p^{(n,R)}(t,\cdot,\cdot):B_{n}(p_{n},R)\times B_{n}(p_{n},R)\rightarrow[0,\infty)$ and $\phi^{(R)}_{n,j}:B_{n}(p_{n},R)\rightarrow\mathbb{R}$ are Hölder continuous. To be precise, there exist constants $C$ and $\Theta$, such that
  1. (1)
    for all $n\geq 1$ and $x_{1},x_{2},y_{1},y_{2}\in B_{n}(p_{n},R)$,
    $\left|p^{(n,R)}(t,x_{1},y_{1})-p^{(n,R)}(t,x_{2},y_{2})\right|\leq\frac{C}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}d_{n}((x_{1},y_{1}),(x_{2},y_{2}))^{\Theta}.$ (4.5)
  2. (2)
    for all $n\geq 1$ and $x,y\in B_{n}(p_{n},R)$,
    $\left|\phi^{(R)}_{n,j}(x)-\phi^{(R)}_{n,j}(y)\right|\leq C\frac{V_{n}(R)^{1/2}}{V_{n}(R/j)}\left({\frac{j}{R}}\right)^{\Theta}d_{n}(x,y)^{\Theta}.$ (4.6)
Proof.
  1. (1)
    According to [GT12, Corollary 5.12], there exist constants $\Theta$ and $C_{H}$ that depends only on the constants appeared in two sided heat kernel estimates, such that for all $(x_{1},y_{1})$ and $(x_{2},y_{2})$ belongs to $B_{n}(p,R)\times B_{n}(p,R)$, if $d_{n}((x_{1},y_{1}),(x_{2},y_{2}))\leq\Psi_{n}^{-1}(t)$, there holds
    $\left|p^{(n,R)}(t,x_{1},y_{1})-p^{(n,R)}(t,x_{2},y_{2})\right|$
    $\leq$ $\left|p^{(n,R)}(t,x_{1},y_{1})-p^{(n,R)}(t,x_{1},y_{2})\right|+\left|p^{(n,R)}(t,x_{1},y_{2})-p^{(n,R)}(t,x_{2},y_{2})\right|$
    $\lesssim$ $\frac{1}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}\left(d_{n}(x_{1},x_{2})^{\Theta}+d_{n}(y_{1},y_{2})^{\Theta}\right)$
    $\lesssim$ $\frac{1}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}d_{n}((x_{1},y_{1}),(x_{2},y_{2}))^{\Theta}.$
    We also have, by the on-diagonal estimates of heat kernel, that if $d_{n}((x_{1},y_{1}),(x_{2},y_{2}))\geq\Psi_{n}^{-1}(t)$, then
    $\left|p^{(n,R)}(t,x_{1},y_{1})-p^{(n,R)}(t,x_{2},y_{2})\right|$ $\lesssim\frac{1}{V_{n}(\Psi_{n}^{-1}(t))}=\frac{1}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}\cdot(\Psi_{n}^{-1}(t))^{\Theta}$ (4.7)
    $\lesssim\frac{1}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}d_{n}((x_{1},y_{1}),(x_{2},y_{2}))^{\Theta}.$ (4.8)
    Combining these two inequalities, we see that for all $n\geq 1$ and $x_{1},x_{2},y_{1},y_{2}\in B_{n}(p_{n},R)$, the estimate (4.5) holds.
  2. (2)
    Using (4.5) and the fact that $\phi_{n,j}$ is also a eigenfunction of the operator $P^{(n,R)}_{t}$ with eigenvalue $e^{-\lambda_{n,j}t}$, we have that for all $t>0$,
    $\left|\phi_{n,j}^{(R)}(x)-\phi_{n,j}^{(R)}(y)\right|$ $=\left|e^{\lambda_{n,j}^{(R)}t}\int_{B_{n}(p_{n},R)}\left(p^{(n,R)}(t,x,z)-p^{(n,R)}(t,y,z)\right)\phi_{n,j}^{(R)}(z)\mathop{}\!\mathrm{d}m_{n}(z)\right|$
    $\lesssim\frac{e^{\lambda_{n,j}^{(R)}t}}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}d_{n}(x,y)^{\Theta}\left\lVert\phi_{n,j}^{(R)}\right\rVert_{L^{2}\left(B_{n}(p_{n},R)\right)}m_{n}(B_{n}(p_{n},R))^{1/2}$
    $\lesssim\frac{e^{\lambda_{n,j}^{(R)}t}}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}d_{n}(x,y)^{\Theta}V_{n}(R)^{1/2}$
    (4.3)${\lesssim}\frac{e^{\frac{C}{\Psi_{n}(R/j)}t}}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}d_{n}(x,y)^{\Theta}V_{n}(R)^{1/2}$
    By choosing $t=\Psi_{n}(R/j)$, we complete the proof of (4.6).
5 Heat kernel on limit space
In this section, we construct the heat kernel on the balls of limit spaces. We first need the following Arzelà–Ascoli type theorem for (possibly) discontinuous functions.
Lemma 5.1 ([Kig23, Lemma D.1]).
Let $(X,d)$ be a totally bounded metric space. Let $u_{n}:X\rightarrow\mathbb{R}$ for any $n\in\mathbb{N}$. Assume that there exists a nondecreasing function $\eta:[0,\infty)\rightarrow[0,\infty)$ and a sequence $\{\delta_{n}\}_{n\in\mathbb{N}}$ of non-negative numbers such that $\lim_{t\downarrow 0}\eta(t)=0$, $\lim_{n\rightarrow\infty}\delta_{n}=0$, $\sup_{n\in\mathbb{N},x\in X}|u_{n}(x)|<\infty$ and
$|u_{n}(x)-u_{n}(y)|\leq\eta(d(x,y))+\delta_{n}\quad\text{for all }x,y\in X\text{ and }n\in\mathbb{N}.$
Then there exists a subsequence $\{u_{n_{k}}\}_{k\in\mathbb{N}}$ and $u\in C(X)$ with
$|u(x)-u(y)|\leq\eta(d(x,y))\quad\text{for all }x,y\in X,$
such that $\sup_{x\in X}|u_{n_{k}}(x)-u(x)|\rightarrow 0$ as $k\rightarrow\infty$.
Definition 5.2.
For any fixed $n\geq 1$, $R>0$, define the family of approximate heat kernels as functions on $(0,\infty)\times B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$ by
$q^{(n,R)}(t,x,y):=p^{(n,R)}(t,g_{n}(x),g_{n}(y)),\quad x,y\in B_{\infty}(p_{\infty},R),\ t>0.$
In the same way, we define the family of approximate $j$-th eigenfunctions as functions on $B_{\infty}(p_{\infty},R)$ by
$\varphi_{n,j}^{(R)}(x):=\phi_{n,j}^{(R)}(g_{n}(x)),\quad x\in B_{\infty}(p_{\infty},R).$
Proposition 5.3.
Fix $R>0$. There exist positive numbers $\{\lambda^{R}_{j}\}_{j=1}^{\infty}$, Hölder continuous functions $\{\varphi^{(R)}_{j}\}_{j=1}^{\infty}$ on $B_{\infty}(p_{\infty},R)$, Hölder continuous functions $\{\widetilde{q}^{(R)}(t,\cdot,\cdot)\}_{t\in\mathbb{Q}_{+}}$ on $B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$, and a common subsequence $\{n_{k}\}_{k\in\mathbb{N}}$ such that
  1. (1)
    for each $t\in\mathbb{Q}_{+}$,
    $\sup_{x,y\in B_{\infty}(p_{\infty},R)}|\widetilde{q}^{(R)}(t,x,y)-q^{(n_{k},R)}(t,x,y)|\rightarrow 0\text{ as }k\rightarrow\infty.$ (5.1)
  2. (2)
    for each $j\geq 1$,
    $|\lambda^{(R)}_{n_{k},j}-\lambda^{R}_{j}|\rightarrow 0\text{ as }k\rightarrow\infty\text{ and }\lambda^{(R)}_{j}\lesssim\frac{1}{\Psi_{\infty}(R/j)}.$ (5.2)
  3. (3)
    for each $j\geq 1$,
    $\sup_{x\in B_{\infty}(p_{\infty},R)}|\varphi_{j}^{R}(x)-\varphi^{R}_{n_{k},j}(x)|$ $\rightarrow 0\text{ as }k\rightarrow\infty$ (5.3)
    $\left\lVert\varphi_{j}^{(R)}\right\rVert_{L^{\infty}\left(B_{n}(p_{n},R)\right)}$ $\lesssim\inf_{t\in(0,\infty)}\frac{e^{\lambda_{j}^{(R)}t}}{V_{\infty}(\Psi_{\infty}^{-1}(t))^{1/2}}.$
  4. (4)
    for $\text{for all }t\in\mathbb{Q}_{+},\ x,y\in B_{\infty}(p_{\infty},R)$ and $\delta>0$,
    the series $\text{the series }\widetilde{q}^{(R)}(t,x,y)=\sum_{j=1}^{\infty}\exp(-\lambda_{j}^{(R)}t)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y)\quad$ (5.4)
    converges uniformly on $\left(\mathbb{Q}\cap[\delta,\infty)\right)\times B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$
Proof.
  1. (1)
    We first construct $q^{(R)}$. By Proposition 4.6, for each fixed $t>0$,
    $\phantom{\leq\ }\left|q^{(n,R)}(t,x_{1},y_{1})-q^{(n,R)}(t,x_{2},y_{2})\right|$ (5.5)
    $=\left|p^{(n,R)}(t,g_{n}(x_{1}),g_{n}(y_{1}))-p^{(n,R)}(t,g_{n}(x_{2}),g_{n}(y_{2}))\right|$ (5.6)
    $\lesssim\frac{1}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}d_{n}((g_{n}(x_{1}),g_{n}(y_{1})),(g_{n}(x_{2}),g_{n}(y_{2})))^{\Theta}$ (5.7)
    $\lesssim\frac{1}{(\Psi_{n}^{-1}(t))^{\Theta}V_{n}(\Psi_{n}^{-1}(t))}\left(d_{\infty}((x_{1},y_{1}),(x_{2},y_{2}))+2\epsilon_{n}\right)^{\Theta}$ (5.8)
    $\lesssim\frac{1}{t^{\Theta/\beta^{\prime}}V_{l}(t^{1/\beta^{\prime}})}\left(d_{\infty}((x_{1},y_{1}),(x_{2},y_{2}))^{\Theta}+\epsilon^{\Theta}_{n}\right),$ (5.9)
    combine which and the uniform heat kernel bounds, we can apply Lemma 5.1 with
    $\eta(s)=\frac{C}{t^{\Theta/\beta^{\prime}}V_{l}(t^{1/\beta^{\prime}})}s^{\Theta}\text{ and }\delta_{n}=\frac{C}{t^{\Theta/\beta^{\prime}}V_{l}(t^{1/\beta^{\prime}})}\epsilon^{\Theta}_{n},$
    and conclude that there is a continuous function, denoted by $q^{(R)}(t,x,y)$, defined on $B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$ such that, along a subsequence (which may depend on $t$ and $R$),
    $\sup_{x,y\in B_{\infty}(p_{\infty},R)}|q^{(R)}(t,x,y)-p^{(n,R)}(t,g_{n}(x),g_{n}(y))|\rightarrow 0\text{ as }n\rightarrow\infty.$
    By a standard diagonal argument, we can find a common subsequence such that the convergence holds for all $t\in\mathbb{Q}_{+}$ along that subsequence.
  2. (2)
    By Proposition 4.2, the sequence $\{\lambda^{(R)}_{n,j}\}_{n=1}^{\infty}$ is bounded for each $j$, we can find a subsequence and a $\lambda^{(R)}_{j}$ such that $\lambda^{(R)}_{n,j}\rightarrow\lambda^{(R)}_{j}$ as $n\rightarrow\infty$, which have the given upper and lower bound.
  3. (3)
    By (4.4), (4.6) and using the same argument as (1), we can apply Lemma 5.1 and find continuous functions $\varphi^{(R)}_{j}$ on $B_{\infty}(p_{\infty},R)$ such that along some subsequence (which may depends on $j$),
    $\sup_{x\in B_{\infty}(p_{\infty},R)}|\varphi_{j}^{R}(x)-\phi^{R}_{n,j}(g_{n}(x))|\rightarrow 0\text{ as }n\rightarrow\infty.$
    By a standard diagonal argument again, we can find a common subsequence $\{n_{k}\}$ such that the convergence in (5.1), (5.2) and (5.3) holds along this subsequence for all $j\geq 1$ and all $t\in\mathbb{Q}_{+}$.
  4. (4)
    From (4.1) we know that for $t\in\mathbb{Q}_{+}$,
    ${q}^{(n_{k},R)}(t,x,y)=\sum_{j=1}^{\infty}\exp(-\lambda_{n_{k},j}^{(R)}t)\varphi_{n_{k},j}^{(R)}(x)\varphi_{n_{k},j}^{(R)}(y)\quad\text{for all }x,y\in B_{\infty}(p_{\infty},R).$
    By (5.1), (5.2), (5.3) and the uniform convergence of (4.1), we can take the limit on both side and conclude that the series in (LABEL:e.ration4) convergent. Fix $\delta>0$. Let us show that the series (LABEL:e.ration4) converges uniformly on $\left(\mathbb{Q}\cap[\delta,\infty)\right)\times B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$. In fact, by (5.3),
    $\left\lVert\varphi_{j}^{(R)}\right\rVert_{L^{\infty}(B_{\infty}(p_{\infty},R))}\lesssim\frac{\exp({\lambda_{j}^{(R)}\delta/4})}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/4))^{1/2}}.$ (5.10)
    Therefore for $\delta\leq t\in\mathbb{Q}_{+}$,
    $\left|\exp(-\lambda_{j}^{(R)}t)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y)\right|$ $\lesssim\exp(-\lambda_{j}^{(R)}\delta)\frac{\exp({\lambda_{j}^{(R)}\delta/2})}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/4))}$ (5.11)
    $\lesssim\frac{\exp(-{\lambda_{j}^{(R)}\delta/2})}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/4))}$ (5.12)
    By Mercer’s theorem on trace [Dav07, Proposition 5.6.9],
    $\sum_{j=1}^{\infty}\exp(-\lambda_{j}^{(R)}\delta/2)$ $=\sum_{j=1}^{\infty}\left(\lim_{k\rightarrow\infty}\exp(-\lambda_{n_{k},j}^{(R)}\delta/2)\right)$
[…]
    Therefore the Weierstrass M-test is applicable to show the uniform convergence in (LABEL:e.ration4).
Now we extend the convergence in (5.1) and the expansion in (LABEL:e.ration4) from $t\in\mathbb{Q}_{+}$ to all $t>0$.
Proposition 5.4.
Fix $R>0$.
  1. (1)
    For $x,y\in B_{\infty}(p_{\infty},R)$ and $t>0$, the limit $\lim_{\mathbb{Q}_{+}\ni s\rightarrow t}\widetilde{q}^{(R)}(s,x,y)$ exists and
    ${q}^{(R)}(t,x,y):=\lim_{\mathbb{Q}_{+}\ni s\rightarrow t}\widetilde{q}^{(R)}(s,x,y)=\sum_{j=1}^{\infty}\exp(-\lambda_{j}^{(R)}t)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y),$ (5.13)
    and the series in (5.13) convergent uniformly on $(t,x,y)\in[\delta,\infty)\times B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$ for any $\delta>0$.
  2. (2)
    The function $q^{(R)}(\cdot,\cdot,\cdot):(0,\infty)\times B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)\to\mathbb{R}$ is continuous. For each $t>0$, $q^{(R)}(t,\cdot,\cdot)$ is Hölder continuous on $B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$.
  3. (3)
    Along the same subsequence in Proposition 5.3, there holds
    $\sup_{x,y\in B_{\infty}(p_{\infty},R)}|{q}^{(R)}(t,x,y)-q^{(n_{k},R)}(t,x,y)|\rightarrow 0\text{ as }k\rightarrow\infty\quad\text{for all }t>0,$ (5.14)
Proof.
  1. (1)
    Let $\delta>0$ and $(t,x,y)\in[\delta,\infty)\times B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)$. By (5.12), the series $\sum_{j=1}^{\infty}\exp(-\lambda_{j}^{(R)}s)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y)$ is dominated by the convergent series $C\sum_{j=1}^{\infty}\frac{\exp(-\lambda_{j}^{(R)}\delta/2)}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/4))}$. Thus we can apply dominated convergence theorem and conclude that
    ${q}^{(R)}(t,x,y)$ $:=\lim_{\mathbb{Q}_{+}\ni s\rightarrow t}\widetilde{q}^{(R)}(s,x,y)$
    $=\lim_{\mathbb{Q}_{+}\ni s\rightarrow t}\sum_{j=1}^{\infty}\exp(-\lambda_{j}^{(R)}s)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y)=\sum_{j=1}^{\infty}\exp(-\lambda_{j}^{(R)}t)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y).$
    which gives (5.13).
  2. (2)
    The continuity of $q^{(R)}(\cdot,\cdot,\cdot)$ follows from the uniform convergent property of (5.13). The Hölder continuity of $q^{(R)}(t,\cdot,\cdot)$ follows from the Hölder continuity of $\tilde{q}^{(R)}(t,\cdot,\cdot)$ in Proposition 5.3 and (5.26) below.
  3. (3)
    Let $\delta\leq s\leq t$.
    |$\phantom{=\ }\absolutevalue{\left(\exp(-\lambda_{j}^{(R)}s)-\exp(-\lambda_{j}^{(R)}t)\right)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y)}$ | (5.15)
    $=\left(\exp\left(-\lambda_{j}^{(R)}\left(s-\frac{1}{2}\delta\right)\right)-\exp\left(-\lambda_{j}^{(R)}\left(t-\frac{1}{2}\delta\right)\right)\right)\absolutevalue{\exp\left(-\lambda_{j}^{(R)}\delta/2\right)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y)}$ | (5.16)
    (5.12)${\lesssim}\frac{\exp(-{\lambda_{j}^{(R)}\delta/4})}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/8))}\left(\exp\left(-\lambda_{j}^{(R)}\left(s-\frac{1}{2}\delta\right)\right)-\exp\left(-\lambda_{j}^{(R)}\left(t-\frac{1}{2}\delta\right)\right)\right)$ ) (5.17)
    $=\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/8))}\left(\exp\left(-\lambda_{j}^{(R)}\left(s-\frac{1}{4}\delta\right)\right)-\exp\left(-\lambda_{j}^{(R)}\left(t-\frac{1}{4}\delta\right)\right)\right)$ (5.18)
    By definition of $q^{(R)}$ and Mercer’s theorem [Dav07, Proposition 5.6.9],
    $\phantom{=\ }\left|{q}^{(R)}(s,x,y)-{q}^{(R)}(t,x,y)\right|$
    $=\left|\sum_{j=1}^{\infty}\left(\exp(-\lambda_{j}^{(R)}s)-\exp(-\lambda_{j}^{(R)}t)\right)\varphi_{j}^{(R)}(x)\varphi_{j}^{(R)}(y)\right|$
    (5.18)${\lesssim}\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/8))}\sum_{j=1}^{\infty}\left(\exp\left(-\lambda_{j}^{(R)}\left(s-\frac{1}{4}\delta\right)\right)-\exp\left(-\lambda_{j}^{(R)}\left(t-\frac{1}{4}\delta\right)\right)\right)$ ) (5.19)
    $\leq\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/8))}\liminf_{k\to\infty}\sum_{j=1}^{\infty}\left(\exp\left(-\lambda_{n_{k},j}^{(R)}\left(s-\frac{1}{4}\delta\right)\right)-\exp\left(-\lambda_{n_{k},j}^{(R)}\left(t-\frac{1}{4}\delta\right)\right)\right)$ (5.20)
    $=\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(\delta/8))}\liminf_{k\to\infty}\left(\int_{B_{n}(p_{n},R)}\left(p^{(n_{k},R)}\left(s-\frac{1}{4}\delta,x,x\right)-p^{(n_{k},R)}\left(t-\frac{1}{4}\delta,x,x\right)\right)\mathop{}\!\mathrm{d}m_{n}(x)\right)$ (5.21)
    By [GT12, Corollary 5.7],
    |$\phantom{=\ }\absolutevalue{\int_{B_{n}(p_{n},R)}\left(p^{(n,R)}\left(s-\frac{1}{4}\delta,x,x\right)-p^{(n,R)}\left(t-\frac{1}{4}\delta,x,x\right)\right)\mathop{}\!\mathrm{d}m_{n}(x)}$ | (5.22)
    $\leq\int_{B_{n}(p_{n},R)}\left(\int_{t-\frac{1}{4}\delta}^{s-\frac{1}{4}\delta}\absolutevalue{\frac{\partial}{\partial r}p^{(n,R)}\left(r,x,x\right)}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}m_{n}(x)$ (5.23)
    $\lesssim\int_{B_{n}(p_{n},R)}\left(\int_{t-\frac{1}{4}\delta}^{s-\frac{1}{4}\delta}\frac{2}{r}\frac{1}{V_{n}(\Psi^{-1}_{n}(r))}\mathop{}\!\mathrm{d}r\right)\mathop{}\!\mathrm{d}m_{n}(x)$ (5.24)
    $\lesssim\frac{8}{3\delta}\frac{V_{n}(R)}{V_{n}(\Psi^{-1}_{n}(3\delta/4))}\absolutevalue{s-t}$ | (5.25)
    Combine (5.21) and (5.25), we see that
    $\left|{q}^{(R)}(s,x,y)-{q}^{(R)}(t,x,y)\right|\lesssim\frac{1}{\delta}\frac{V_{\infty}(R)}{V_{\infty}(\Psi^{-1}_{\infty}(\delta/8))^{2}}\absolutevalue{s-t}.$ | . (5.26)
    Similarly by (5.25),
    $\left|q^{(n_{k},R)}(t,x,y)-q^{(n_{k},R)}(t,x,y)\right|\lesssim\frac{1}{\delta}\frac{V_{n}(R)}{V_{n_{k}}(\Psi^{-1}_{n_{k}}(\delta/8))^{2}}\absolutevalue{s-t}.$ | . (5.27)
    For each fixed $\delta>0$ and any $t>\delta>0$, if $t\in(0,\infty)\setminus\mathbb{Q}$, we choose $s\in(\delta,t)\cap\mathbb{Q}$, and by (5.26) and (5.27),
    $\left|{q}^{(R)}(t,x,y)-q^{(n_{k},R)}(t,x,y)\right|$
    $\leq$ $\left|{q}^{(R)}(t,x,y)-{q}^{(R)}(s,x,y)\right|+\left|{q}^{(R)}(s,x,y)-q^{(n_{k},R)}(s,x,y)\right|+\left|q^{(n_{k},R)}(s,x,y)-q^{(n_{k},R)}(t,x,y)\right|$
    $\lesssim$ $\frac{1}{\delta}\left(\frac{V_{n_{k}}(R)}{V_{n_{k}}(\Psi^{-1}_{n_{k}}(\delta/8))^{2}}+\frac{V_{\infty}(R)}{V_{\infty}(\Psi^{-1}_{\infty}(\delta/8))^{2}}\right)(t-s)+\left|{q}^{(R)}(s,x,y)-q^{(n_{k},R)}(s,x,y)\right|,$
    which implies
    $\phantom{\leq\ }\lim_{k\to\infty}\sup_{x,y\in B_{\infty}(p_{\infty},R)}\left|{q}^{(R)}(t,x,y)-q^{(n_{k},R)}(t,x,y)\right|$ (5.28)
    $\lesssim\frac{1}{\delta}\frac{V_{\infty}(R)}{V_{\infty}(\Psi^{-1}_{\infty}(\delta/8))^{2}}(t-s)+\lim_{k\to\infty}\sup_{x,y\in B_{\infty}(p_{\infty},R)}\left|{q}^{(R)}(s,x,y)-q^{(n_{k},R)}(s,x,y)\right|$ (5.29)
    (LABEL:e.ration4)${\lesssim}\frac{1}{\delta}\frac{V_{\infty}(R)}{V_{\infty}(\Psi^{-1}_{\infty}(\delta/8))^{2}}(t-s)$ ) (5.30)
    for any $s\in(\delta,t)\cap\mathbb{Q}$. Letting $\mathbb{Q}\ni s\to t$, we obtain (5.14).
6 Dirichlet form on limit space
In this section, we will show that the function $q^{(R)}$ given in Proposition 5.4 gives a heat kernel and a semigroup on $B_{\infty}(p_{\infty},R)$. By letting $R\to\infty$, we will obtain a heat kernel on $X_{\infty}$.
Proposition 6.1.
For each $R>0$, the functions $\{\varphi_{j}^{(R)}\}_{j=1}^{\infty}$ are orthonormal in $L^{2}(B_{\infty}(p_{\infty},R),m_{\infty})$.
Proof.
Let $1\leq j,k<\infty$. By weak convergence of measures and the Hölder continuity of $\phi_{j}^{(R)}$ in Proposition 5.3,
$\phantom{=\ \ }\int_{B_{\infty}(p_{\infty},R)}\varphi_{j}^{(R)}(x)\varphi_{k}^{(R)}(x)\mathop{}\!\mathrm{d}m_{\infty}(x)$
$=\lim_{n\rightarrow\infty}\int_{B_{\infty}(p_{\infty},R)}\varphi_{j}^{(R)}(x)\varphi_{k}^{(R)}(x)\mathop{}\!\mathrm{d}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}(p_{n},R_{n})}}\right)$
=(3.1)${=}\lim_{n\rightarrow\infty}\int_{\overline{B_{n}(p_{n},R_{n})}}{\mathds{1}}_{B_{\infty}(p_{\infty},R)}(f_{n}(x))\varphi_{j}^{(R)}(f_{n}(x))\varphi_{k}^{(R)}(f_{n}(x))\mathop{}\!\mathrm{d}m_{n}(x)$ )
$:=\lim_{n\rightarrow\infty}(I_{1,n}+I_{2,n})$
where
$I_{1,n}$ $:=\int_{B_{n}(p_{n},R)}\varphi_{j}^{(R)}(f_{n}(x))\varphi_{k}^{(R)}(f_{n}(x))\mathop{}\!\mathrm{d}m_{n}(x)$
$I_{2,n}$ $:=\int_{X_{n}}\left({\mathds{1}}_{A_{n}}(x)-{\mathds{1}}_{C_{n}}(x)\right)\varphi_{j}^{(R)}(f_{n}(x))\varphi_{k}^{(R)}(f_{n}(x))\mathop{}\!\mathrm{d}m_{n}(x)$
$A_{n}$ $:=\left(\overline{B_{n}(p_{n},R_{n})}\cap f_{n}^{-1}\left(B_{\infty}(p_{\infty},R)\right)\right)\setminus B_{n}(p_{n},R)\subset B_{n}(p_{n},R+2\epsilon_{n})\setminus B_{n}(p_{n},R),$
$C_{n}$ $:=B_{n}(p_{n},R)\setminus\left(\overline{B_{n}(p_{n},R_{n})}\cap f_{n}^{-1}\left(B_{\infty}(p_{\infty},R)\right)\right)\subset B_{n}(p_{n},R)\setminus B_{n}(p_{n},R-2\epsilon_{n}),$
since $\left(\overline{B_{n}(p_{n},R_{n})}\cap f_{n}^{-1}\left(B_{\infty}(p_{\infty},R)\right)\right)=(B_{n}(p_{n},R)\setminus C_{n})\cup A_{n}$. As a consequence, by Lemma 3.1 we have that
$|I_{2,n}|\leq m_{n}\left(B_{n}(p_{n},R+2\epsilon_{n})\setminus B_{n}(p_{n},R-2\epsilon_{n})\right)\cdot\lVert\varphi_{j}^{R}\rVert_{L^{\infty}}\lVert\varphi_{k}^{R}\rVert_{L^{\infty}}\rightarrow 0\text{ as }n\rightarrow\infty.$
Therefore,
$\int_{B_{\infty}(p_{\infty},R)}\varphi_{j}^{(R)}(x)\varphi_{k}^{(R)}(x)\mathop{}\!\mathrm{d}m_{\infty}(x)=\lim_{n\rightarrow\infty}I_{1,n}$
$=$ $\lim_{n\rightarrow\infty}\int_{B_{n}(p_{n},R)}\varphi_{j}^{(R)}(f_{n}(x))\varphi_{k}^{(R)}(f_{n}(x))-\varphi_{n,j}^{(R)}(f_{n}(x))\varphi_{n,k}^{(R)}(f_{n}(x))\mathop{}\!\mathrm{d}m_{n}(x)$
$\ +\lim_{n\rightarrow\infty}\int_{B_{n}(p_{n},R)}\varphi_{n,j}^{(R)}(f_{n}(x))\varphi_{n,k}^{(R)}(f_{n}(x))-\phi_{n,j}^{(R)}(x)\phi_{n,k}^{(R)}(x)\mathop{}\!\mathrm{d}m_{n}(x)$
$\ +\lim_{n\rightarrow\infty}\int_{B_{n}(p_{n},R)}\phi_{n,j}^{(R)}(x)\phi_{n,k}^{(R)}(x)\mathop{}\!\mathrm{d}m_{n}(x)$
The first term is zero as $(\varphi_{n,j}^{(R)}\circ g_{n})(\varphi_{n,k}^{(R)}\circ g_{n})\rightarrow\varphi_{j}^{(R)}\varphi_{j}^{(R)}$ uniformly by Proposition 5.3 (3), and $m_{n}(B_{n}(p,R))$ is uniformly bounded. The second term is zero, as $\{\phi_{n,j}^{(R)}\phi_{n,k}^{(R)}\}_{1\leq n<\infty}$ is equi-continuous by (4.6), $\varphi_{n,j}^{(R)}\circ f_{n}=\phi_{n,j}^{(R)}\circ(g_{n}\circ f_{n})$ and $d_{n}((g_{n}\circ f_{n})(x),x)\leq\epsilon_{n}\rightarrow 0$ as $n\rightarrow\infty$. The third term is $\delta_{jk}$ since the functions $\{\varphi_{n,j}^{(R)}\}_{j=1}^{\infty}$ are orthonormal in $L^{2}(B_{n}(p_{n},R),m_{n})$. ∎
Proposition 6.2.
$q^{(R)}$ is a heat kernel on $B_{\infty}(p_{\infty},R)$. To be precise,
  1. (1)
    Measurability: The function $q^{(R)}:(0,\infty)\times B_{\infty}(p_{\infty},R)\times B_{\infty}(p_{\infty},R)\to\mathbb{R}$ is measurable with respect to the Borel $\sigma$-algebra.
  2. (2)
    Symmetry: for any $x,y\in B_{\infty}(p_{\infty},R)$, $t>0$, $q^{(R)}(t,x,y)=q^{(R)}(t,y,x)\geq 0$
  3. (3)
    Semigroup property: for all $t,s>0$ and $x,y\in B_{\infty}(p_{\infty},R)$,
    $q^{(R)}(t+s,x,y)=\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,x,z)q^{(R)}(s,z,y)\mathop{}\!\mathrm{d}m_{\infty}(z).$ (6.1)
  4. (4)
    Contraction property: for all $f\in L^{2}(B_{\infty}(p_{\infty},R))$,
    $\left\lVert\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,\cdot,y)f(y)\mathop{}\!\mathrm{d}m_{\infty}(y)\right\rVert_{L^{2}(B_{\infty}(p_{\infty},R))}\leq\left\lVert f\right\rVert_{L^{2}(B_{\infty}(p_{\infty},R))}.$ (6.2)
Proof.
  1. (1)
    Since $q^{(R)}$ is continuous by Proposition 5.4-(2), it is measurable.
  2. (2)
    This follows directly from (5.14).
  3. (3)
    By Proposition 5.4-(1), Proposition 6.1 and Fubini’s theorem,
    $\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,x,z)q^{(R)}(s,z,y)\mathop{}\!\mathrm{d}m_{\infty}(z)$
    =\ = $\int_{B_{\infty}(p_{\infty},R)}\sum_{j,k=1}^{\infty}\exp(-\lambda_{k}^{(R)}t-\lambda_{j}^{(R)}s)\varphi_{k}^{(R)}(x)\varphi_{k}^{(R)}(z)\varphi_{j}^{(R)}(z)\varphi_{j}^{(R)}(y)\mathop{}\!\mathrm{d}m_{\infty}(z)$
    =\ = $\sum_{j,k=1}^{\infty}\exp(-\lambda_{k}^{(R)}t-\lambda_{j}^{(R)}s)\left(\int_{B_{\infty}(p_{\infty},R)}\varphi_{k}^{(R)}(z)\varphi_{j}^{(R)}(z)\mathop{}\!\mathrm{d}m_{\infty}(z)\right)\varphi_{k}^{(R)}(x)\varphi_{j}^{(R)}(y)$
    =\ = $\sum_{j,k=1}^{\infty}\exp(-\lambda_{k}^{(R)}t-\lambda_{j}^{(R)}s)\varphi_{k}^{(R)}(x)\varphi_{j}^{(R)}(y)\delta_{jk}$
    =\ = $\sum_{j}^{\infty}\exp(-\lambda_{j}^{(R)}(t+s))\varphi_{k}^{(R)}(x)\varphi_{j}^{(R)}(y)=q^{(R)}(t+s,x,y).$
  4. (4)
    For any $f\in L^{2}(B_{\infty}(p_{\infty},R))$, by Fubini’s theorem we have
    $\int_{B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,x,y)f(y)\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
    $\leq$ $\ \int_{B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,x,y)\left|f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
    ${=}$ $\ \int_{B_{\infty}(p_{\infty},R)}\left(\sum_{j=1}^{\infty}e^{-\lambda^{(R)}_{j}t}\left\langle\phi_{j}^{(R)},\left|f\right|\right\rangle_{L^{2}(B_{\infty}(p_{\infty},R))}\phi_{j}^{(R)}(x)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
[…]
    which gives (6.2).
We set $q^{(R)}(t,x,y)$ to zero if $x$ or $y$ outside $B_{\infty}(p_{\infty},R)$. By the monotonicity of $p^{(n,R)}$ with respect to $R$ for each $n$, we can see that $0\leq q^{(R_{1})}(t,x,y)\leq q^{(R_{2})}(t,x,y)$ for all $t>0$, $0<R_{1}\leq R_{2}$ and $x,y\in X_{\infty}$, so that the limit
$q(t,x,y):=\lim_{R\rightarrow\infty}q^{(R)}(t,x,y)=\lim_{\mathbb{Q}_{+}\ni R\rightarrow\infty}q^{(R)}(t,x,y)=\lim_{j\rightarrow\infty}q^{(R_{j})}(t,x,y).$ (6.3)
is monotone and thus exists. We first prove that function $q(t,x,y)$ is conservative.
Proposition 6.3.
The function $q$ is conservative, in the sense that:
$\int_{X_{\infty}}q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}=1,\text{ for any $x\in X_{\infty}$, $t>0$.}$ (6.4)
Proof.
Fix $x\in X_{\infty}$ and choose $R$ large enough so that $g_{n}(x)\in B_{n}(p_{n},R/3)$ for all $n$ sufficiently large, say $R>4d_{\infty}(p_{\infty},x)+1$. As $q^{(R)}(t,x,\cdot)$ is continuous on $B_{\infty}(p_{\infty},R)$, we can use the property of weak convergence and estimates the measure of annulus as the proof in Proposition 6.1 and show that
$\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}=$ $\lim_{n\rightarrow\infty}\int_{B_{n}(p_{n},R)}q^{(R)}(t,x,f_{n}(y))\mathop{}\!\mathrm{d}m_{n}(y)$ (6.5)
$=$ $\lim_{n\rightarrow\infty}\int_{B_{n}(p_{n},R)}q^{(R)}(t,x,f_{n}(y))-p^{(n,R)}(t,g_{n}(x),g_{n}(f_{n}(y)))\mathop{}\!\mathrm{d}m_{n}(y)$ (6.6)
$\ +\lim_{n\rightarrow\infty}\int_{B_{n}(p_{n},R)}p^{(n,R)}(t,g_{n}(x),g_{n}(f_{n}(y)))-p^{(n,R)}(t,g_{n}(x),y)\mathop{}\!\mathrm{d}m_{n}(y)$ (6.7)
$\ +\lim_{n\rightarrow\infty}\int_{B_{n}(p_{n},R)}p^{(n,R)}(t,g_{n}(x),y)\mathop{}\!\mathrm{d}m_{n}(y).$ (6.8)
The first two terms vanish by (5.14) and the same argument as Proposition 6.1. The rest is to estimate the third term. By [GT12, Corollary 3.20], we see that for some $c,\gamma\in(0,\infty)$,
$1\geq\int_{B_{n}(p_{n},R)}p^{(n,R)}(t,g_{n}(x),y)\mathop{}\!\mathrm{d}m_{n}(y)$ $=\mathbb{P}^{g_{n}(x)}(t<\tau_{B_{n}(p_{n},R)})$ (6.9)
(since $g_{n}(x)\in B_{n}(p_{n},R/3)$) $\geq\mathbb{P}^{g_{n}(x)}(t<\tau_{B_{n}(g_{n}(x),R/2)})$ (6.10)
$=1-\mathbb{P}^{g_{n}(x)}(\tau_{B_{n}(g_{n}(x),R/2)}\leq t)$ (6.11)
$\geq 1-c\exp\left(-\gamma\left({\frac{\Psi_{n}(R)}{t}}\right)^{1/(\beta^{\prime}-1)}\right)$ (6.12)
Let $n\to\infty$, we see from (6.8) that for any $R>4d_{\infty}(p_{\infty},x)+1$,
$1\geq\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(y)\geq 1-c\exp\left(-\gamma\left({\frac{\Psi_{\infty}(R)}{t}}\right)^{1/(\beta^{\prime}-1)}\right).$ (6.13)
Letting $R\to\infty$, we obtain (6.4) since $\Psi_{\infty}(R)\to\infty$. ∎
Lemma 6.4.
Assume (A2) and (A3). Let $\Phi_{n}:=\Phi_{\Psi_{n}}$, $1\leq n\leq\infty$ be defined by (1.3). Then
$\Phi_{n}\to\Phi_{\infty}\text{ converges uniformly on compact sets.}$.
Proof.
Since $\Psi_{n}(1)\to\Psi_{\infty}(1)$ as $n\to\infty$, there is a positive number $C>0$ such that $C^{-1}\leq\Psi_{n}(1)\leq C$ for all $1\leq n\leq\infty$ and therefore
[…]
Fix $0<m<M$. We first show that $\Phi_{n}\to\Phi_{\infty}$ uniformly on $[m,M]$. We may assume that $M>C$. For any $s\in[m,M]$,
$\Phi_{n}(s)\geq\sup_{r>0}\left({\frac{m}{r}-\frac{1}{(C_{0}C)^{-1}(r^{\beta}\wedge r^{\beta^{\prime}})}}\right):=C_{2}=C_{2}(m,C,C_{0},\beta,\beta^{\prime}).$
Let $C_{3}:=C_{3}(m,M,C,C_{0},\beta,\beta^{\prime})=MC_{2}^{-1}$, we have for all $s\in[m,M]$ that
$\sup_{r>C_{3}}\left({\frac{s}{r}-\frac{1}{\Psi_{n}(r)}}\right)\leq\frac{M}{r}<C_{0}C(\beta-1)\left({\frac{m(C_{0}C)^{-1}}{\beta}}\right)^{\beta/(\beta-1)}\leq\Phi_{n}(s)=\sup_{r>0}\left({\frac{s}{r}-\frac{1}{\Psi_{n}(r)}}\right).$ (6.14)
Let $C_{4}:=C_{4}(C_{0},M,\beta,\beta^{\prime})=(C_{0}CM)^{-1/(\beta-1)}\vee(C_{0}CM)^{-1/(\beta^{\prime}-1)}$, we have
$\sup_{0<r<C_{4}}\left(\frac{s}{r}-\frac{1}{\Psi_{n}(r)}\right)\leq\frac{M}{r}-\frac{1}{(C_{0}C)^{-1}(r^{\beta}\vee r^{\beta^{\prime}})}<0.$ (6.15)
Hence (6.14) and (6.15) gives
[…]
Since $\Psi_{n}\to\Psi_{\infty}$ uniformly on $[C_{4},C_{3}]$, we conclude that $\Phi_{n}\to\Phi_{\infty}$ uniformly on $[m,M]$, for every $0<m<M$. Next, we show that $\Phi_{n}\to\Phi_{\infty}$ uniformly on $[0,M]$, for every $M>0$. In fact, by [Mur20, Lemma 2.10], for any $\epsilon>0$, we can find a $m>0$ small enough, independent of $n$, such that |$\absolutevalue{\Phi_{n}(s)}\lesssim\epsilon$ for all $s\in[0,m]$ and $1\leq n\leq\infty$. Then choose $N$ large enough such that |$\absolutevalue{\Phi_{n}(s)-\Phi_{\infty}(s)}\lesssim\epsilon$ for all $s\in[m,M]$ and all $n\geq N$. Therefore, the uniform convergence holds on $[0,M]$ for all $M>0$. ∎
Theorem 6.5.
Let $q(t,x,y)$ be defined in (6.3).
  1. (1)
    Measurability: for any $t>0$, $q$ is a measureable function on $(0,\infty)\times X_{\infty}\times X_{\infty}$.
  2. (2)
    Markovian and conservative property: for any $x,y\in X_{\infty}$, $t>0$, $q(t,x,y)\geq 0$ and
    $\int_{X_{\infty}}q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}=1.$ (6.16)
  3. (3)
    Symmetry: for any $x,y\in X_{\infty}$, $t>0$, $q(t,x,y)=q(t,y,x)$.
  4. (4)
    Semigroup property: for all $t,s>0$,
    $q(t+s,x,y)=\int_{X_{\infty}}q(t,x,z)q(s,z,y)\mathop{}\!\mathrm{d}m_{\infty}(z)$ (6.17)
  5. (5)
    Heat kerneal bounds: $q$ satisfies $\mathrm{HKE}(\Psi_{\infty})$.
Proof.
  1. (1)
    Since $q^{(R)}$ is measurable by Proposition 5.4-(1), $q$ is also measurable as it is the pointwise limit of $q^{(R)}$.
  2. (2)
    The positivity of $q(t,\cdot,\cdot)$ is obvious. That $q$ being conservative have been proved in Proposition 6.3.
  3. (3)
    The symmetry of $q$ follows from the symmetry of $q^{(R)}$ in Proposition (1).
  4. (4)
    Taking $R\rightarrow\infty$ on both sides of (6.1), and using the monotone convergence theorem, we have the semigroup property (6.17) of $q$.
  5. (5)
    By the upper bound of $p^{(n)}$,
    $q^{(n,R)}(t,x,y)=p^{(n,R)}(t,g_{n}(x),g_{n}(y))$ $\leq p^{(n)}(t,g_{n}(x),g_{n}(y))$
    $\lesssim\frac{C_{1}}{V_{n}(\Psi_{n}^{-1}(t))}\exp\left(-c_{2}t\Phi_{n}\left(c_{3}\frac{d_{n}(g_{n}(x),g_{n}(y))}{t}\right)\right)$
    $\leq\frac{C_{1}}{V_{n}(\Psi_{n}^{-1}(t))}\exp\left(-c_{2}t\Phi_{n}\left(c_{3}\frac{d_{\infty}^{\beta}(x,y)-\epsilon_{n}}{t}\right)\right)$
    Letting $n\rightarrow\infty$, by Lemma 6.4 we see that
    $q^{(R)}(t,x,y)\lesssim\frac{C_{1}}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\exp\left(-c_{2}t\Phi_{\infty}\left(c_{3}\frac{d_{\infty}^{\beta}(x,y)}{t}\right)\right)\text{ for all }R>1.$
    Let $R\rightarrow\infty$ and use (6.3) we obtain the upper bound of $q$.
    To show the lower bound, let $0<\delta<\delta^{\prime}$ be given later. Let $t>0$ and $x,y\in X_{\infty}$ with $d_{\infty}(x,y)\leq\delta\Psi_{\infty}^{-1}(t)$. Therefore $d_{\infty}(x,y)<\delta^{\prime}\Psi_{\infty}^{-1}(t)$ and we may find $N=N(t,x,y)$ large enough such that if $n>N$ then $d_{n}(g_{n}(x),g_{n}(y))<\delta^{\prime}\Psi_{n}^{-1}(t)$. Recall that under our assumption, $\mathrm{HKE}(\Psi_{n})$ is quantitively equivalent to the local lower estimate $\mathrm{LLE}(\Psi_{n})$ [BGK12, Theorem 3.1], so there exist constants $\epsilon\in(0,1)$ and $c>0$ such that
    $p^{(B_{n}(g_{n}(x),\epsilon\Psi_{n}^{-1}(t))}(t,g_{n}(x),g_{n}(y))\geq\frac{c}{V_{n}(\Psi_{n}^{-1}(t))}$
    for all $t>0$ and all $d_{n}(g_{n}(x),g_{n}(y))<\epsilon\Psi_{n}^{-1}(t)$. We choose $\delta^{\prime}=\epsilon$ and $\delta=\epsilon/2$. For $R$ large enough such that $R>d_{\infty}(p_{\infty},x)+2\delta^{\prime}\Psi_{\infty}^{-1}(t)+1$, we have by triangle inequality that $g_{n}(y)\in B_{n}(g_{n}(x),\delta^{\prime}\Psi_{n}^{-1}(t))\subset B_{n}(p_{n},R_{n})$. By the monotonicity of $p^{(n,R)}$ with respect to $R$, we have
    $q^{(n,R)}(t,x,y)=p^{(n,R)}(t,g_{n}(x),g_{n}(y))\geq p^{(B_{n}(g_{n}(x),\epsilon\Psi_{n}^{-1}(t))}(t,g_{n}(x),g_{n}(y))\geq\frac{c}{V_{n}(\Psi_{n}^{-1}(t))}.$
    Using Proposition 5.3 and letting $n\to\infty$ along some subsequence, we have $q^{(R)}(t,x,y)\geq\frac{c}{V_{\infty}(\Psi_{\infty}^{-1}(t))}$. Letting $R\to\infty$, we have
    $q(t,x,y)\geq\frac{c}{V_{\infty}(\Psi_{\infty}^{-1}(t))}.$
Theorem 6.6.
The operators $\{Q_{t}\}_{t>0}$ defined by
$Q_{t}f(x):=\int_{X_{\infty}}q(t,x,y)f(y)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.18)
give a Markovian strongly continuous semigroup of bounded operators, in the sense of the following:
  1. (1)
    Each $Q_{t}$ is a bounded symmetric operator with domain ${L^{2}(X_{\infty},m_{\infty})}$.
  2. (2)
    Semigroup property: $Q_{t}Q_{s}=Q_{t+s}$, $t,s>0$.
  3. (3)
    Contraction property: $\lVert Q_{t}f\rVert_{L^{2}(X_{\infty},m_{\infty})}\leq\lVert f\rVert_{L^{2}(X_{\infty},m_{\infty})}$, $t>0$, $f\in L^{2}(X_{\infty},m_{\infty})$.
  4. (4)
    Strongly continuity: $\lVert Q_{t}f-f\rVert_{L^{2}(X_{\infty},m_{\infty})}\rightarrow 0$ as $t\rightarrow 0$, for all $f\in L^{2}(X_{\infty},m_{\infty})$
  5. (5)
    Conservativity: $Q_{t}\mathds{1}_{X_{\infty}}=\mathds{1}_{X_{\infty}}$ for any $t>0$.
  6. (6)
    Markovian property: $0\leq Q_{t}f\leq 1$ $m_{\infty}$-a.e. whenever $f\in L^{2}(X_{\infty},m_{\infty})$, $0\leq f\leq 1$ $m_{\infty}$-a.e.
Proof.
  1. (1)
    This follows directly from the symmetry of $q$. The boundedness follows from the proof of contraction property, see (3) below.
  2. (2)
    The semigroup property follows from (6.17).
  3. (3)
    For any $f\in L^{2}(X_{\infty},m_{\infty})$, by the monotone convergence theorem,
    $\lVert Q_{t}f\rVert^{2}_{L^{2}(X_{\infty},m_{\infty})}$ $=\int_{X_{\infty}}\left(\int_{X_{\infty}}q(t,x,y)f(y)\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
    $\leq\int_{X_{\infty}}\left(\int_{X_{\infty}}q(t,x,y)\left|f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
    $=\lim_{R\rightarrow\infty}\int_{B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)}q^{(R)}(t,x,y)\left|f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
    (6.2)${\leq}\lVert f\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}.$
  4. (4)
    We first show that $\lVert Q_{t}f-f\rVert_{L^{2}(X_{\infty},m_{\infty})}\rightarrow 0$ as $t\rightarrow 0$ holds for all $f\in C_{c}(X_{\infty})$. We may assume that $\mathrm{supp}(f)\subset B_{\infty}(p_{\infty},R/2)\subset B_{\infty}(p_{\infty},R)$. Since $f$ is uniformly continuous, for each $\epsilon>0$, there exists a $\delta>0$ such that $\left|f(x)-f(y)\right|<\epsilon$ whenever $d_{\infty}(x,y)\leq\delta$. By the conservative of $q$,
    $\int_{X_{\infty}}\left|Q_{t}f(x)-f(x)\right|^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
[…]
    where
    $I_{1}(t)$ $:=2\int_{B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)\cap B_{\infty}(x,\delta)}q(t,x,y)\left|f(x)-f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x),$
    $I_{2}(t)$ $:=2\int_{B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)\setminus B_{\infty}(x,\delta)}q(t,x,y)\left|f(x)-f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x),$
    $I_{3}(t)$ $:=2\int_{B_{\infty}(p_{\infty},R)}\left(\int_{X_{\infty}\setminus B_{\infty}(p_{\infty},R)}q(t,x,y)\left|f(x)-f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x),$
    $I_{4}(t)$ $:=\int_{X_{\infty}\setminus B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)}q(t,x,y)\left|f(x)-f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x).$
    By the uniform continuity of $f$ and Proposition 6.3,
    $I_{1}(t)\leq 2\epsilon^{2}\int_{B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)\cap B_{\infty}(x,\delta)}q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)\lesssim V_{\infty}(R)\epsilon^{2}.$ (6.19)
    By [GT12, Lemma 3.19], $q(t,x,y)\lesssim(V_{\infty}(\Psi_{\infty}^{-1}(t)))^{-1}\exp(-c\Psi_{\infty}(\delta)^{1/(\beta^{\prime}-1)}t^{-1/(\beta^{\prime}-1)})$ if $d_{\infty}(x,y)\geq\delta$ and $t$ is small, we have
    $I_{2}(t)$ $\lesssim V_{\infty}(R)^{3}(V_{\infty}(\Psi_{\infty}^{-1}(t)))^{-2}\exp(-2c\Psi_{\infty}(\delta)^{1/(\beta^{\prime}-1)}t^{-1/(\beta^{\prime}-1)})\lVert f\rVert^{2}_{L^{\infty}(X_{\infty},m_{\infty})}$ (6.20)
    $\lesssim V_{\infty}(R)\left(\frac{R}{\Psi_{\infty}^{-1}(t)}\right)^{2\alpha^{\prime}}\exp(-2c\Psi_{\infty}(\delta)^{1/(\beta^{\prime}-1)}t^{-1/(\beta^{\prime}-1)})\lVert f\rVert^{2}_{L^{\infty}(X_{\infty},m_{\infty})}$
    $\lesssim V_{\infty}(R)\left(\frac{\Psi_{\infty}(R)}{t}\right)^{2\alpha^{\prime}/\beta}\exp(-2c\Psi_{\infty}(\delta)^{1/(\beta^{\prime}-1)}t^{-1/(\beta^{\prime}-1)})\lVert f\rVert^{2}_{L^{\infty}(X_{\infty},m_{\infty})}\to 0\ (t\to 0^{+}).$
    By [Mur20, Lemma 2.10], we have $t\Phi_{\infty}(s/t)\gtrsim(s^{\beta^{\prime}}/t)^{1/(\beta^{\prime}-1)}$ whenever $s\geq t$. Therefore, if $t\leq\Psi_{\infty}(R/2)$,
    $\phantom{\lesssim}\int_{X_{\infty}\setminus B_{\infty}(x,R)}q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.21)
    $\lesssim\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\int_{B_{\infty}(x,R/2)^{c}}\exp\left({-ct\Phi_{\infty}\left({\frac{d_{\infty}(x,y)}{t}}\right)}\right)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.22)
    $\lesssim\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\int_{B_{\infty}(x,R/2)^{c}}\exp\left(-c\left(\frac{d_{\infty}(x,y)^{\beta^{\prime}}}{t}\right)^{1/(\beta^{\prime}-1)}\right)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.23)
    $=\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\sum_{j=0}^{\infty}\int_{B_{\infty}(x,2^{j}R)\setminus B_{\infty}(x,2^{j-1}R)}\exp\left(-c\left(\frac{d_{\infty}(x,y)^{\beta^{\prime}}}{t}\right)^{1/(\beta^{\prime}-1)}\right)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.24)
    $\lesssim\sum_{j=0}^{\infty}\left(\frac{2^{j-1}R}{t^{1/\beta^{\prime}}}\right)^{\alpha^{\prime}}\exp\left(-\frac{c}{2}\left(\frac{2^{j-1}R}{t^{1/\beta^{\prime}}}\right)^{\beta^{\prime}/(\beta^{\prime}-1)}\right)\lesssim\exp\left(-\frac{c}{2}\left(\frac{R^{\beta^{\prime}}}{t}\right)^{1/(\beta^{\prime}-1)}\right)$ (6.25)
    Therefore, for $t\leq\Psi_{\infty}(R/2)$
    $I_{3}(t)$ $\lesssim\int_{B_{\infty}(p_{\infty},R)}\left(\int_{X_{\infty}\setminus B_{\infty}(p_{\infty},R)}q(t,x,y)\left|f(x)-f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$ (6.26)
    $=\int_{B_{\infty}(p_{\infty},R/2)}\left|f(x)\right|^{2}\left(\int_{X_{\infty}\setminus B_{\infty}(p_{\infty},R)}q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$ (6.27)
    $\lesssim\int_{B_{\infty}(p_{\infty},R/2)}\left|f(x)\right|^{2}\left(\int_{B_{\infty}(x,R/2)^{c}}q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$ (6.28)
    (6.25)${\lesssim}\lVert f\rVert^{2}_{L^{2}(X_{\infty},m_{\infty})}\exp\left(-2c\left(\frac{R^{\beta^{\prime}}}{t}\right)^{1/(\beta^{\prime}-1)}\right)\to 0\ (t\to 0^{+})$ ) (6.29)
    By Minkowski’s inequality,
    $I_{4}(t)$ $\lesssim\int_{X_{\infty}\setminus B_{\infty}(p_{\infty},R)}\left(\int_{B_{\infty}(p_{\infty},R)}q(t,x,y)\left|f(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
    $\lesssim\left(\int_{B_{\infty}(p_{\infty},R)}\left|f(y)\right|\left(\int_{X_{\infty}\setminus B_{\infty}(p_{\infty},R)}q(t,x,y)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)\right)^{1/2}\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}$
    $\lesssim\left(\int_{B_{\infty}(p_{\infty},R/2)}\left|f(y)\right|\left(\int_{X_{\infty}\setminus B_{\infty}(y,R/2)}q(t,x,y)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)\right)^{1/2}\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}$
    (6.25)${\lesssim}\frac{\lVert f\rVert^{2}_{L^{2}(X_{\infty},m_{\infty})}}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\exp\left(-2c\left(\frac{R^{\beta^{\prime}}}{t}\right)^{1/(\beta^{\prime}-1)}\right)$ ) (6.30)
    $\lesssim\frac{\lVert f\rVert^{2}_{L^{2}(X_{\infty},m_{\infty})}}{V_{\infty}(R)}\left({\frac{\Psi_{\infty}(R)}{t}}\right)^{\alpha^{\prime}/\beta}\exp\left(-2c\left(\frac{R^{\beta^{\prime}}}{t}\right)^{1/(\beta^{\prime}-1)}\right)\to 0\ (t\to 0^{+}).$ (6.31)
    Combining (6.19), (6.20), (6.29), (6.31) and letting $t\rightarrow 0$ we see that
    $\limsup_{t\rightarrow 0}\lVert Q_{t}f-f\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}\lesssim V_{\infty}(R)\epsilon^{2}.$
    Since $\epsilon$ is arbitrary, we conclude that $\lim_{t\rightarrow 0}\lVert Q_{t}f-f\rVert_{L^{2}(X_{\infty},m_{\infty})}=0$.
    For general $f\in L^{2}(X_{\infty},m_{\infty})$, given any $\epsilon>0$, we can find $g\in C_{c}(X_{\infty})$ such that $\lVert f-g\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}\leq\epsilon$. By the contraction property, $\lVert Q_{t}f-Q_{t}g\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}\leq\lVert f-g\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}\leq\epsilon$. Thus by trangle inequality, $\limsup_{t\rightarrow 0}\lVert Q_{t}f-f\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}\leq 2\epsilon+\limsup_{t\rightarrow 0}\lVert Q_{t}g-g\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}=2\epsilon$. Again, since $\epsilon$ is arbitrary, we conclude that $\lim_{t\rightarrow 0}\lVert Q_{t}f-f\rVert_{L^{2}(X_{\infty},m_{\infty})}=0$.
  5. (5)
    This follows from Proposition 6.3.
  6. (6)
    This follows immediately from the fact that $q\geq 0$ and the conservative property.
Now we can prove the existence of the Dirichlet form on the limit space:
Proof of Theorem 1.2-(2).
  1. (i)
    By the general theory of Dirichlet forms [FOT11, Chapter 1], this Markovian strongly continuous semigroup $\{Q_{t}\}_{t>0}$ corresponds to a unique Dirichlet form, which is denoted by $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$.
  2. (ii)
    $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ is symmetric and conservative since $\{Q_{t}\}$ is. The heat kernel estimates is given in Theorem 6.5-(5).
  3. (iii)
    Since $\{Q_{t}\}$ is conservative and admits two-sided heat kernel estimates $\mathrm{HKE}(\Psi_{\infty})$ by Theorem 6.5-(5), we can apply [Lie15, Proposition 3.2] and see that the set $\mathcal{C}:=\{Q_{t}f:f\in C_{0}(X_{\infty})\cap L^{2}(X_{\infty},m_{\infty}),t>0\}$ is a core of $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$. In fact, by [Lie15, Proposition 3.2] and [FOT11, Lemma 1.3.3 (i) and (iii)], we see that $\mathcal{C}\subset C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$, $\mathcal{C}$ is dense in $C_{0}(X_{\infty})$ in the uniform norm and dense in $C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ in $\mathcal{E}_{\infty,1}$. We now prove that $C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ is dense in $\mathcal{F}_{\infty}$ in the $\mathcal{E}_{\infty,1}$ norm, so $\mathcal{C}$ is also dense in $\mathcal{F}_{\infty}$ in the $\mathcal{E}_{\infty,1}$ norm. By [FOT11, Lemma 1.3.3 (iii)], it suffices to prove that $Q_{t}(L^{2}(X_{\infty},m_{\infty}))\subset C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ for each $t>0$.
    From [FOT11, Lemma 1.3.3 (i)], we know that $Q_{t}(L^{2}(X_{\infty},m_{\infty}))\subset\mathcal{F}$. Fix $\epsilon>0$, $f\in L^{2}(X_{\infty},m_{\infty})$ and let $R>0$ large enough such that $\lVert f\rVert_{L^{2}(X_{\infty}\setminus B_{\infty}(p_{\infty},R),m_{\infty})}\leq\epsilon$, then
    |$\absolutevalue{Q_{t}f(z)}$ | $\leq\int_{B_{\infty}(p_{\infty},R)}\absolutevalue{f(y)}q(t,z,y)\mathop{}\!\mathrm{d}m_{\infty}(y)+\int_{X_{\infty}\setminus B_{\infty}(p_{\infty},R)}\absolutevalue{f(y)}q(t,z,y)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.32)
    $\leq\int_{B_{\infty}(p_{\infty},R)}\absolutevalue{f(y)}q(t,z,y)\mathop{}\!\mathrm{d}m_{\infty}(y)+\epsilon q(2t,z,z)^{1/2}$ (6.33)
    The first term can be made small if $d_{\infty}(p_{\infty},z)$ is large enough by $\mathrm{HKE}(\Psi_{\infty})$, so $Q_{t}f\in C_{0}(X_{\infty})$ for each $t>0$. This completes the proof that $\mathcal{C}$ is a core.
  4. (iv)
    By Theorem 6.6-(5) and [Lie15, Proposition 3.2], $\{Q_{t}\}$ or $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ corresponds to a conservative Hunt process on $X_{\infty}$ with continuous path, therefore $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ is strongly local by [FOT11, Theorem 4.5.4].
7 Mosco convergence of Dirichlet forms
In this section, we will show that along some subsequence, the Dirichlet form $\left(\mathcal{E}_{\infty},\mathcal{F}_{\infty}\right)$ is the Mosco limit of $\left(\mathcal{E}_{n},\mathcal{F}_{n}\right)$ as $n\rightarrow\infty$. For each $1\leq n\leq\infty$, we identify $(\mathcal{E}_{n},\mathcal{F}_{\infty})$ with the quadratic energy form $\mathcal{E}_{n}$ on $L^{2}(X_{n},m_{n})$ defined by
$\mathcal{E}_{n}(u):=\begin{dcases}\mathcal{E}_{n}(u,u)&\text{if $u\in\mathcal{F}_{n}$,}\\ \infty&\text{if $u\in L^{2}(X_{n},m_{n})\setminus\mathcal{F}_{n}$.}\end{dcases}$ (7.1)
The following definition extends the $L^{2}$ convergence of functions to varying spaces.
Definition 7.1.
Assume that the Hilbert spaces $L^{2}(X_{n},m_{n})$ converge to $L^{2}(X_{\infty},m_{\infty})$ (see Definition 3.3).
  1. (1)
    A sequence $\left\{u_{n}\right\}_{1\leq n<\infty}$ of real functions with $u_{n}\in L^{2}\left(X_{n},m_{n}\right)$ is said to $L^{2}$ converges to a real function $u\in L^{2}\left(X_{\infty},m_{\infty}\right)$, or $u_{n}\to u$ in $L^{2}$, if there exists a sequence of functions $\left\{v_{j}\right\}_{1\leq j<\infty}\subset\mathcal{C}$ converges to $u$ in $L^{2}\left(X_{\infty},m_{\infty}\right)$ such that
    $\lim_{j\rightarrow\infty}\limsup_{n\rightarrow\infty}\left\|\upphi_{n}v_{j}-u_{n}\right\|_{L^{2}\left(X_{n},m_{n}\right)}=0.$
  2. (2)
    A sequence $\left\{u_{n}\right\}_{1\leq n<\infty}$ of real functions with $u_{n}\in L^{2}\left(X_{n},m_{n}\right)$ is said to $L^{2}$ weakly converges to a real function $u\in L^{2}\left({X}_{\infty},m_{\infty}\right)$ if
    $\lim_{n\rightarrow\infty}\left\langle u_{n},w_{n}\right\rangle_{L^{2}\left(X_{n},m_{n}\right)}=\langle u,w\rangle_{L^{2}\left(X_{\infty},m_{\infty}\right)}$
    for any sequence $\left\{w_{n}\right\}$ converges to $w$.
  3. (3)
    Let $A_{n}$ be a bounded linear operator on $L^{2}(X_{n},m_{n})$, $1\leq n\leq\infty$. We say $A_{n}\to A_{\infty}$ in strong operator topology if $A_{n}u_{n}\to A_{\infty}u_{\infty}$ in $L^{2}$ whenever $u_{n}\to u_{\infty}$ in $L^{2}$.
The following definition of Mosco convergence was originally due to Mosco [Mos94, Definition 2.1.1] for a fixed space and developed to varying spaces by Kuwae and Shioya [KS03, Definition 2.11].
Definition 7.2 (Mosco Convergence).
We say that quadratic forms $\mathcal{E}_{n}$ on $L^{2}\left(X_{n},m_{n}\right)$ Mosco converge to a quadratic form $\mathcal{E}_{\infty}$ on $L^{2}\left(X_{\infty},m_{\infty}\right)$ if the following holds:
  1. (1)
    If a sequence $\left\{u_{n}\right\}_{1\leq n<\infty}$ of real functions with $u_{n}\in L^{2}\left(X_{n},m_{n}\right)$ is $L^{2}$ weakly converge to $u_{\infty}\in L^{2}\left(X_{\infty},m_{\infty}\right)$, then
    $\mathcal{E}_{\infty}\left(u_{\infty}\right)\leq\liminf_{n\rightarrow\infty}\mathcal{E}_{n}\left(u_{n}\right)$
  2. (2)
    For any $u$ in $L^{2}\left(X_{\infty},m_{\infty}\right)$, there exists $\left\{u_{n}\right\}_{1\leq n<\infty}$ of real functions with $u_{n}\in L^{2}\left(X_{n},m_{n}\right)$ which $L^{2}$ converges to $u$ and
    $\mathcal{E}_{\infty}\left(u\right)=\lim_{n\rightarrow\infty}\mathcal{E}_{n}\left(u_{n}\right)$
The following theorem of Mosco [Mos94, Corollary 2.6.1] and Kuwae-Shioya [KS03, Theorem 2.4] is used to characterize Mosco convergence in the convergence of semigroups.
Theorem 7.3 (Mosco, Kuwae-Shioya).
Let $\mathcal{E}_{n}$ be a quadratic form on $L^{2}(X_{n},m_{n})$ for each $1\leq n\leq\infty$. Let $\{P_{t}^{(n)}\}_{t>0}$ be their corresponding semigroups. Then $\mathcal{E}_{n}$ Mosco converges to $\mathcal{E}_{\infty}$ if and only if $P_{t}^{(n)}$ converges to $P_{t}^{(\infty)}$ in the strong operator topology, for each $t>0$.
Proof of Theorem 1.2-(3).
Since we take only countable many subsequences on Section 6 and in (6.3), we can find a common subsequence such that the estimates in Proposition 5.4 holds for all $R\in\mathbb{Q}_{+}\cup\{R_{j}\}_{j=1}^{\infty}$. We will show that along this subsequence, $P_{t}^{(n)}u_{n}\rightarrow Q_{t}u$ in $L^{2}$ for $u_{n}\rightarrow u$ in $L^{2}$ for every $t>0$, which gives Theorem 1.2-(3) by Theorem 7.3.
By definition, there exist some $v_{j}\in C_{c}\left(X_{\infty},d_{\infty}\right)$ such that
$v_{j}\rightarrow u\text{ in }L^{2}\left(X_{\infty},m_{\infty}\right)$ (7.2)
and
$\lim_{j\rightarrow\infty}\limsup_{n\rightarrow\infty}\left\|\upphi_{n}v_{j}-u_{n}\right\|_{L^{2}\left(X_{n},m_{n}\right)}=0.$ (7.3)
where $\upphi_{n}$ is defined in (3.4). We may assume that the sequences in Proposition 2.3 have the property that $\{\epsilon_{n}\}\subset(0,1/8)$ and $\{R_{j}\}\subset[1,\infty)$. Moreover we may assume that $\mathrm{supp}\left(v_{j}\right)\subset B_{\infty}\left(p_{\infty},R_{j}-1\right)\subset B_{\infty}\left(p_{\infty},R_{j}\right)$ so that $\mathrm{supp}(\upphi_{n}v_{j})\subset B_{n}(p_{n},R_{j})$. Let $\phi_{j}\in C_{c}\left(B_{\infty}\left(p_{\infty},R_{j}\right)\right)$ be a cut-off function such that
$0\leq\phi_{j}\leq 1$ and $\left.\phi_{j}\right|_{B_{\infty}\left(p_{\infty},R_{j}-R_{j}^{-\delta}\right)}=1$,
where the constant $\delta>0$ will be chosen later. Define
$w_{j}(x):=\phi_{j}(x)\cdot Q_{t}^{\left(R_{j}\right)}v_{j}(x),\quad x\in X_{\infty}.$ (7.4)
Then $w_{j}\in C_{c}\left(X_{\infty},d_{\infty}\right)$. We claim that:
$w_{j}\rightarrow Q_{t}u\text{ in }L^{2}\left(X_{\infty},m_{\infty}\right)$ (7.5)
and
$\lim_{j\rightarrow\infty}\limsup_{n\rightarrow\infty}\left\|\upphi_{n}w_{j}-P_{t}^{(n)}u_{n}\right\|_{L^{2}\left(X_{n},m_{n}\right)}=0.$ (7.6)
which will imply, by Definition 7.1, that $P_{t}^{(n)}u_{n}\rightarrow Q_{t}u$ in $L^{2}$.
  1. (i)
    Proof of (7.5).
    By triangle inequality,
    $\left\|w_{j}-Q_{t}u\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}=\left\|\phi_{j}\cdot Q_{t}^{\left(R_{j}\right)}v_{j}-Q_{t}u\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}\leq\sum_{k=1}^{4}I_{k}(j)$ (7.7)
    where
    $I_{1}(j)$ $:=\left\|\phi_{j}\cdot Q_{t}^{\left(R_{j}\right)}v_{j}-Q_{t}^{\left(R_{j}\right)}v_{j}\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}$ (7.8)
    $I_{2}(j)$ $:=\left\|Q_{t}^{\left(R_{j}\right)}v_{j}-Q_{t}^{\left(R_{j}\right)}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}$ (7.9)
    $I_{3}(j)$ $:=\left\|Q_{t}^{\left(R_{j}\right)}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)-Q_{t}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}$ (7.10)
    $I_{4}(j)$ $:=\left\|Q_{t}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)-Q_{t}u\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}.$ (7.11)
    Recall that we set $q^{(R)}(t,x,y)$ to be zero if either $x$ or $y$ is outside $B_{\infty}\left(p_{\infty},R\right)$. Since $V_{u}$ is doubling, we can find a constant $\delta>0$ such that
    $R^{-(\delta+1)\gamma}V_{u}(R)\to 0\ \text{ as }R\to\infty,$ (7.12)
    where $\gamma$ is the constant in Proposition 3.1.
    1. (a)
      By a direct estimate,
      $I_{1}(j)$ $\leq\sup_{x\in B_{\infty}\left(p_{\infty},R_{j}\right)}\left|Q_{t}^{\left(R_{j}\right)}v_{j}(x)\right|\cdot m_{\infty}\left(B_{\infty}\left(p_{\infty},R_{j}\right)\backslash B_{\infty}\left(p_{\infty},R_{j}-R_{j}^{-\delta}\right)\right)$
      $\lesssim R_{j}^{-(\delta+1)\gamma}V_{\infty}(R_{j})\sup_{x\in B_{\infty}\left(p_{\infty},R_{j}\right)}\left|\int_{B_{\infty}\left(p_{\infty},R_{j}\right)}\left(q^{\left(R_{j}\right)}(t,x,y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(y)\right|^{1/2}\left\|v_{j}\right\|_{L^{2}\left(B_{\infty}\left(p_{\infty},R_{j}\right)\right)}$
      (by Cauchy-Schwarz’s inequality and Proposition 3.1)
      $\lesssim R_{j}^{-(\delta+1)\gamma}V_{\infty}(R_{j})\sup_{x\in B_{\infty}\left(p_{\infty},R_{j}\right)}q^{\left(R_{j}\right)}(2t,x,x)\lVert u\rVert_{L^{2}\left(X_{\infty},m_{\infty}\right)}$
      (by (7.2), the symmetry and semigroup property in Proposition 6.2)
[…]
    2. (b)
      For the second term, by the contraction property (6.2) of $Q^{(R_{j})}$ and triangle inequality,
      $I_{2}(j)$ $\leq\left\|v_{j}-u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}$
      $\leq\left\|v_{j}-u\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}+\left\|u-u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}\rightarrow 0\quad(j\rightarrow\infty).$ (7.14)
    3. (c)
      By definition, for any $x\in X_{\infty}$, whenever $R_{j}\geq 4d_{\infty}(p_{\infty},x)+1$, we have
      $\phantom{\leq\ }\left|Q_{t}^{\left(R_{j}\right)}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)(x)-Q_{t}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)(x)\right|^{2}$ (7.15)
      $\leq\left(\int_{X_{\infty}}\left|q(t,x,y)-q^{\left(R_{j}\right)}(t,x,y)\right|\left|u(y)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}$ (7.16)
      $\leq\left(\int_{X_{\infty}}\left|q(t,x,y)-q^{\left(R_{j}\right)}(t,x,y)\right|^{2}\mathop{}\!\mathrm{d}m_{\infty}(y)\right)\lVert u\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}$ (7.17)
      $\leq\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\left(\int_{X_{\infty}}q(t,x,y)-q^{\left(R_{j}\right)}(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(y)\right)\lVert u\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}$ (7.18)
      (6.13)${\lesssim}\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\exp\left(-\gamma\left({\frac{\Psi_{\infty}(R_{j})}{t}}\right)^{1/(\beta^{\prime}-1)}\right)\lVert u\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}$ (7.19)
      $\rightarrow 0\quad(j\to\infty).$ (7.20)
      Therefore the function $x\mapsto\left|Q_{t}^{\left(R_{j}\right)}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)(x)-Q_{t}\left(u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right)(x)\right|$ goes to 00 pointwisely. Meanwhile, they are dominated by $2Q_{t}(|u|)$, which is in $L^{2}\left(X_{\infty},m_{\infty}\right)$, so we may apply the dominated convergence theorem and conclude that $I_{3}(j)\rightarrow 0$ as $j\rightarrow\infty$.
    4. (d)
      By the contractive property of $Q_{t}$ in Theorem 6.6, we have
      $I_{4}(j)\leq\left\|u-u{\mathds{1}}_{B_{\infty}\left(p_{\infty},R_{j}\right)}\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}\rightarrow 0\quad(j\rightarrow\infty).$ (7.21)
    Combining the estimates (7.13), (7.14), (7.20), (7.21) and (7.7), we complete the proof of (7.5).
  2. (ii)
    Proof of (7.6).
    By triangle inequality,
    $\phantom{\leq\ }\left\|\upphi_{n}w_{j}-P_{t}^{(n)}u_{n}\right\|_{L^{2}\left(X_{n},m_{n}\right)}$ (7.22)
    $\leq\left\|\upphi_{n}w_{j}-P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)\right\|_{L^{2}\left(X_{n},m_{n}\right)}+\left\|P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)-P_{t}^{(n)}u_{n}\right\|_{L^{2}\left(X_{n},m_{n}\right)}$
    $\leq\sum_{k=1}^{4}I_{k}(n,j),$ (7.23)
    where
    $I_{1}(n,j)$ $:=\left\|\upphi_{n}w_{j}-\upphi_{n}\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)\right\|_{L^{2}\left(X_{n},m_{n}\right)}$ (7.24)
    $I_{2}(n,j)$ $:=\left\|\upphi_{n}\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)-P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)\right\|_{L^{2}\left(X_{n},m_{n}\right)}$ (7.25)
    $I_{3}(n,j)$ $:=\left\|P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)-P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)\right\|_{L^{2}\left(X_{n},m_{n}\right)}$ (7.26)
    $I_{4}(n,j)$ $:=\left\|P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)-P_{t}^{(n)}u_{n}\right\|_{L^{2}\left(X_{n},m_{n}\right)}$ (7.27)
    We estimate $I_{k}(n,j)$, $1\leq k\leq 4$ term by term as follows.
    1. (a)
      Estimate $I_{1}(n,j)$.
      By definition of $\upphi_{n}$ and $w_{j}$, if $x\in\overline{B_{n}\left(p_{n},R_{n}\right)}$,
      $\left|\upphi_{n}w_{j}(x)-\upphi_{n}\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)(x)\right|=\left(1-\phi_{j}\left(f_{n}(x)\right)\right)\cdot\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)\left(f_{n}(x)\right)$ (7.28)
      Since $Q_{t}^{\left(R_{j}\right)}v_{j}$ vanishes outside ${B_{\infty}\left(p_{\infty},R_{j}\right)}$ and is Hölder continuous on ${B_{\infty}\left(p_{\infty},R_{j}\right)}$, we conclude that, as $n\rightarrow\infty$,
      $I_{1}(n,j)^{2}=$ $\int_{X_{n}}\left|\upphi_{n}w_{j}(x)-\upphi_{n}\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)(x)\right|^{2}\mathop{}\!\mathrm{d}m_{n}(x)$
      $=$ $\ \int_{\overline{B_{n}\left(p_{n},R_{n}\right)}}\left(1-\phi_{j}\left(f_{n}(x)\right)\right)^{2}\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)^{2}\left(f_{n}(x)\right)\mathrm{d}m_{n}(x)$
      $=$ $\ \int_{B_{\infty}\left(p_{\infty},R_{j}\right)}\left(1-\phi_{j}(y)\right)^{2}\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)^{2}(y)\mathop{}\!\mathrm{d}\left(f_{n}\right)_{\#}\left(m_{n}|_{\overline{B_{n}\left(p_{n},R_{n}\right)}}\right)(y)$
      (${\rightarrow}$ $\ \int_{B_{\infty}\left(p_{\infty},R_{j}\right)}\left(1-\phi_{j}(y)\right)^{2}\left(Q_{t}^{\left(R_{j}\right)}v_{j}\right)^{2}(y)\mathop{}\!\mathrm{d}m_{\infty}(y)=I_{1}(j)^{2},$ (7.29)
      which means
      $\limsup_{n\rightarrow\infty}I_{1}(n,j)=I_{1}(j).$ (7.30)
      Therefore
[…]
    2. (b)
      Estimate $I_{2}(n,j)$.
      For $n$ sufficiently large, since the mass between $B_{n}\left(p_{n},R_{j}+2\epsilon_{n}\right)\setminus B_{n}\left(p_{n},R_{j}\right)$ is small, we have by triangle inequality that
      $I_{2}(n,j)^{2}\lesssim$ $\int_{B_{n}\left(p_{n},R_{j}\right)}\left(Q_{t}^{\left(R_{j}\right)}{v_{j}}\left(f_{n}(x)\right)-P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)(x)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$
      $\lesssim$ $\int_{B_{n}\left(p_{n},R_{j}\right)}\left(Q_{t}^{\left(R_{j}\right)}{v_{j}}\left(f_{n}(x)\right)-P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)\left(g_{n}\left(f_{n}(x)\right)\right)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$
      $+\int_{B_{n}\left(p_{n},R_{j}\right)}\left(P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)\left(g_{n}\left(f_{n}(x)\right)\right)-P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)(x)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$
      $=$ $\int_{f_{n}\left(B_{n}\left(p_{n},R_{j}\right)\right)}\left(Q_{t}^{\left(R_{j}\right)}{v_{j}}(y)-P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)\left(g_{n}(y)\right)\right)^{2}\mathop{}\!\mathrm{d}\left(f_{n}\right)_{\#}m_{n}(y)$
      $+\int_{B_{n}\left(p_{n},R_{j}\right)}\left(P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)\left(g_{n}\left(f_{n}(x)\right)\right)-P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)(x)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$
      ::=:= $I_{2,1}(n,j)+I_{2,2}(n,j)$
      To estimates $I_{2,1}(n,j)$,
      $\left|Q_{t}^{\left(R_{j}\right)}{v_{j}}(y)-P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)\left(g_{n}(y)\right)\right|$
      $=$ $\left|\int_{B_{\infty}\left(p_{\infty},R_{j}\right)}q^{\left(R_{j}\right)}(t,y,z)v_{j}(z)\mathop{}\!\mathrm{d}m_{\infty}(z)-\int_{B_{n}\left(p_{n},R_{j}\right)}p^{\left(n,R_{j}\right)}\left(t,g_{n}(y),z\right)v_{j}\left(f_{n}(z)\right)\mathop{}\!\mathrm{d}m_{n}(z)\right|$
      $\leq$ $\left|\int_{B_{\infty}\left(p_{\infty},R_{j}\right)}q^{\left(R_{j}\right)}(t,y,z)v_{j}(z)\mathop{}\!\mathrm{d}m_{\infty}(z)-\int_{B_{n}\left(p_{n},R_{j}\right)}q^{\left(n,R_{j}\right)}\left(t,y,f_{n}(z)\right)v_{j}\left(f_{n}(z)\right)\mathop{}\!\mathrm{d}m_{n}(z)\right|$
      $+\left|\int_{B_{n}\left(p_{n},R_{j}\right)}\left(p^{\left(n,R_{j}\right)}\left(t,g_{n}(y),g_{n}\left(f_{n}(z)\right)\right)-p^{\left(n,R_{j}\right)}\left(t,g_{n}(y),z\right)\right)v_{j}\left(f_{n}(z)\right)\mathop{}\!\mathrm{d}m_{n}(z)\right|$
      $\rightarrow 0\text{ uniformly on $y$, as $n\rightarrow\infty$}$ (7.32)
      by Proposition 3.2 and the equi-continuity of $p^{\left(n,R_{j}\right)}(t,\cdot,\cdot)$, from which we conclude that $\lim\sup_{n\rightarrow\infty}I_{2,1}(n,j)=0$. It is easy to see that, by the equi-continuity of $p^{\left(n,R_{j}\right)}(t,\cdot,\cdot)$ and dominated convergence theorem, $\lim\sup_{n\rightarrow\infty}I_{2,2}(n,j)=0$. Thus
      $\limsup_{n\rightarrow\infty}I_{2}(n,j)=0.$ (7.33)
    3. (c)
      Estimate $I_{3}(n,j)$.
      We decompose the integration in $I_{3}(n,j)$ into two parts:
      $I_{3,1}(n,j)$ $:=\left\|P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)-P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)\right\|_{L^{2}\left(B_{n}\left(p_{n},R_{j}\right),m_{n}\right)}^{2}$ (7.34)
      and $\text{and }I_{3,1}(n,j)$ $:=\left\|P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)-P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)\right\|_{L^{2}\left(B_{n}\left(p_{n},R_{j}\right)^{c},m_{n}\right)}^{2},$ (7.35)
      so that $I_{3}(n,j)^{2}=I_{3,1}(n,j)+I_{3,2}(n,j)$. Given any $\epsilon>0$, we fix a $j_{\epsilon}\in\mathbb{N}$ such that $\left\|v_{j}-v_{j_{\epsilon}}\right\|_{L^{2}\left(X_{\infty},m_{\infty}\right)}<\epsilon$ for all $j\geq j_{\epsilon}$, as $\left\{v_{j}\right\}$ is a Cauchy sequence in $L^{2}\left(X_{\infty},m_{\infty}\right)$.
      Since $\upphi_{n}v_{j}$ is supported in $B_{n}\left(p_{n},R_{j}\right)$ when $n$ is large enough, we have by triangle inequality that
      $\phantom{\leq\ }I_{3,1}(n,j)=\left\|P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)-P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)\right\|_{L^{2}\left(B_{n}\left(p_{n},R_{j}\right),m_{n}\right)}^{2}$
      $=\int_{B_{n}\left(p_{n},R_{j}\right)}\left(\int_{B_{n}\left(p_{n},R_{j}\right)}\left(p^{(n)}(t,x,y)-p^{(n,R_{j})}(t,x,y)\right)\upphi_{n}v_{j}(y)\mathop{}\!\mathrm{d}m_{n}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$
      $\lesssim\int_{B_{n}\left(p_{n},R_{j}\right)}\left(\int_{B_{n}\left(p_{n},R_{j}\right)}\left(p^{(n)}(t,x,y)-p^{(n,R_{j})}(t,x,y)\right)\absolutevalue{\upphi_{n}v_{j}(y)-\upphi_{n}v_{j_{\epsilon}}(y)}\mathop{}\!\mathrm{d}m_{n}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$
      $\quad+\int_{B_{n}\left(p_{n},R_{j}\right)}\left(\int_{B_{n}\left(p_{n},R_{j}\right)}\left(p^{(n)}(t,x,y)-p^{(n,R_{j})}(t,x,y)\right)\absolutevalue{\upphi_{n}v_{j_{\epsilon}}(y)}\mathop{}\!\mathrm{d}m_{n}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)$ (7.36)
      The first term in (7.36) is less than or equal to
      $\left\lVert P^{(n)}_{t}\absolutevalue{\upphi_{n}v_{j}-\upphi_{n}v_{j_{\epsilon}}}\right\rVert_{L^{2}\left(X_{n},m_{n}\right)}^{2}\leq\lVert\upphi_{n}v_{j}-\upphi_{n}v_{j_{\epsilon}}\rVert_{L^{2}\left(X_{n},m_{n}\right)}^{2}$ (7.37)
      which tends to $\lVert v_{j}-v_{j_{\epsilon}}\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}\leq\epsilon^{2}$ as $n\to\infty$ by the uniform continuity of $v_{j}$ and $v_{j_{\epsilon}}$ and the weak convergence of measures. Since $\upphi_{n}v_{j_{\epsilon}}$ is supported in $B_{n}\left(p_{n},R_{j_{\epsilon}}\right)$, we have by Minkowski inequality that, whenever $R_{j}\geq 8R_{j_{\epsilon}}$, the second term in (7.36) is less than or equal to
      $\phantom{\leq\ }\left(\int_{B_{n}\left(p_{n},R_{j_{\epsilon}}\right)}\left(\int_{B_{n}\left(p_{n},R_{j}\right)}\left(p^{(n)}(t,x,y)-p^{(n,R_{j})}(t,x,y)\right)^{2}\mathop{}\!\mathrm{d}m_{n}(x)\right)^{1/2}\left|\upphi_{n}v_{j_{\epsilon}}(y)\right|\mathop{}\!\mathrm{d}m_{n}(y)\right)^{2}$ (7.38)
      $\lesssim\frac{1}{V_{n}(\Psi_{n}^{-1}(t))}\left(\int_{B_{\infty}\left(p_{\infty},R_{j_{\epsilon}}\right)}\exp\left(-\frac{\gamma}{2}\left({\frac{\Psi_{n}(R_{j}-R_{j_{\epsilon}})}{t}}\right)^{1/(\beta^{\prime}-1)}\right)\left|\upphi_{n}v_{j_{\epsilon}}(y)\right|\mathop{}\!\mathrm{d}m_{n}(y)\right)^{2}$ (7.39)
      $\lesssim\frac{1}{V_{n}(\Psi_{n}^{-1}(t))}V_{n}(R_{j_{\epsilon}})\exp\left(-\frac{\gamma}{2}\left({\frac{\Psi_{n}(R_{j}-R_{j_{\epsilon}})}{t}}\right)^{1/(\beta^{\prime}-1)}\right)\lVert\upphi_{n}v_{j_{\epsilon}}\rVert_{L^{2}(X_{n},m_{n})}^{2}$ (7.40)
      $\to\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(t))}V_{\infty}(R_{j_{\epsilon}})\exp\left(-\frac{\gamma}{2}\left({\frac{\Psi_{\infty}(R_{j}-R_{j_{\epsilon}})}{t}}\right)^{1/(\beta^{\prime}-1)}\right)\lVert v_{j_{\epsilon}}\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}$ (7.41)
      as $n\to\infty$. (7.42)
      where we use the upper bound of $p^{(n)}$ and [GT12, Corollary 3.10] in the first inequality, Cauchy-Schwarz’s inequality in the second inequality, and the uniform continuity of $v_{j_{\epsilon}}$ in the third line. Letting $j\to\infty$, we see that the second term in (7.36) vanishes and therefore
      $\limsup_{j\to\infty}\limsup_{n\to\infty}I_{3,1}(n,j)\lesssim\epsilon^{2}.$ (7.43)
      On the other hand, since $P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)=0$ on $B_{n}\left(p_{n},R_{j}\right)^{c}$, by triangle inequality and Minkowski inequality, whenever $R_{j}\geq 8R_{j_{\epsilon}}$,
      $\phantom{\leq\ }\limsup_{n\rightarrow\infty}I_{3,2}(n,j)=\limsup_{n\rightarrow\infty}\left\|P_{t}^{\left(n,R_{j}\right)}\left(\upphi_{n}v_{j}\right)-P_{t}^{(n)}\left(\upphi_{n}v_{j}\right)\right\|_{L^{2}\left(B_{n}\left(p_{n},R_{j}\right)^{c},m_{n}\right)}^{2}$
      $\lesssim\limsup_{n\rightarrow\infty}\left\|P_{t}^{(n)}\left(\upphi_{n}v_{j}-\upphi_{n}v_{j_{\epsilon}}\right)\right\|_{L^{2}\left(X_{n},m_{n}\right)}^{2}+\limsup_{n\rightarrow\infty}\left\|P_{t}^{(n)}\left(\upphi_{n}v_{j_{\epsilon}}\right)\right\|_{L^{2}\left(B_{n}\left(p_{n},R_{j}\right)^{c},m_{n}\right)}^{2}$
      $\lesssim\epsilon^{2}+\limsup_{n\rightarrow\infty}\int_{B_{n}\left(p_{n},R_{j}\right)^{c}}\left(\int_{B_{n}\left(p_{n},R_{j_{\epsilon}}\right)}p^{(n)}(t,x,y)v_{j_{\epsilon}}\left(f_{n}(y)\right)\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}\mathop{}\!\mathrm{d}m_{\infty}(x)$
      $\lesssim\epsilon^{2}+\limsup_{n\rightarrow\infty}\left(\int_{B_{n}\left(p_{n},R_{j_{\epsilon}}\right)}\left\|p^{(n)}(t,\cdot,y)\right\|_{L^{2}\left(B_{n}\left(y,R_{j}-R_{j_{\epsilon}}\right)^{c},m_{n}\right)}\left|v_{j_{\epsilon}}\left(f_{n}(y)\right)\right|\mathop{}\!\mathrm{d}m_{\infty}(y)\right)^{2}$
      $\lesssim\epsilon^{2}+\frac{V_{\infty}(R_{j_{\epsilon}})}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\exp\left(-c\left({\frac{\Psi_{\infty}(R_{j}-R_{j_{\epsilon}})}{t}}\right)^{1/(\beta^{\prime}-1)}\right)\lVert v_{j_{\epsilon}}\rVert_{L^{2}(X_{\infty},m_{\infty})}^{2}.$ (7.44)
      Combining (7.43) and (7.44), we conclude that
      $\limsup_{j\rightarrow\infty}\limsup_{n\rightarrow\infty}I_{3}(n,j)\lesssim\epsilon^{2}.$
      Since $\epsilon$ is arbitrary, we have
      $\lim_{j\rightarrow\infty}\limsup_{n\rightarrow\infty}I_{3}(n,j)=0.$ (7.45)
    4. (d)
      Estimtae $I_{4}(n,j)$.
      Since $P^{(n)}_{t}$ is contractive, we have $I_{4}(n,j)\leq\left\|\upphi_{n}v_{j}-u_{n}\right\|_{L^{2}\left(X_{n},m_{n}\right)}$. Using (7.3), we have
      $\lim_{j\rightarrow\infty}\limsup_{n\rightarrow\infty}I_{4}(n,j)=0$ (7.46)
    Combining (7.31), (7.33), (7.45), (7.46) and (7.23), we finally have (7.6).
8 Some examples
In this section, we will reconstruct conservative, regular, and strongly local Dirichlet forms on the unbounded Sierpiński carpet and the unbounded Sierpiński gasket in Examples 8.1 and 8.2, respectively, as applications of Theorem 1.2. The primary components of this reconstruction are the reflected Dirichlet forms on pre-carpets and the Brownian motion on Sierpiński gasket cable systems. In Example 8.3, we present an alternative proof of the existence of diffusions with prescribed heat kernel estimates featuring power scaling functions, which constitutes a special case of a recent result by Murugan [Mur24].
Example 8.1.
Let $F_{0}=[0,1]^{2}\subset\mathbb{R}^{2}$ be the unit square, $l=3$, $N=8$ and $S=\{1,\ldots,N\}$. Define $q_{k}$, $k\in S$ by
$q_{1}$ $:=(0,0),$ $q_{2}:=(1/3,0),$ $q_{3}:=(2/3,0),$ $q_{4}:=(2/3,1/3),$
$q_{5}$ $:=(2/3,2/3),$ $q_{6}:=(1/3,2/3),$ $q_{7}:=(0,2/3),$ $q_{8}:=(0,1/3).$
Let $h_{k}:\mathbb{R}^{2}\to\mathbb{R}^{2}$, $k\in S$ denote the similitude $h_{k}(x)=l^{-1}x+q_{k}$. We define a sequence of decreasing compact set $\{F_{n}\}$ inductively by
$F_{n+1}:=\bigcup_{k\in S}h_{k}(F_{n}),\ n\geq 0.$ (8.1)
Let $F_{\infty}:=\bigcap_{n=0}^{\infty}F_{n}$ be the standard planer Sierpiński carpet and let
$X_{0}:=\bigcup_{n\geq 0}l^{n}F_{n}\text{ and }X_{\infty}:=\bigcup_{n\geq 0}l^{n}F_{\infty}$ (8.2)
We call $X_{0}$ the pre-carpet and $X_{\infty}$ the unbounded Sierpiński carpet. Obviously, $X_{\infty}\subsetneq X_{0}$.
  1. (1)
    We denote $d_{\mathbb{R}^{2}}$ be a metric on $\mathbb{R}^{2}$ (and on its subsets) defined by $d_{\mathbb{R}^{2}}((x_{1},y_{1}),(x_{2},y_{2}))=\max(|x_{1}-x_{2}|,|y_{1}-y_{2}|)$. Let $d_{0}$ (resp. $d_{\infty}$) be the intrinsic distance of $X_{0}$ (resp. $X_{\infty}$) with respect to $d_{\mathbb{R}^{2}}$, i.e. $d_{0}(x,y)$ (resp. $d_{\infty}(x,y)$) is the length of the shortest path in $X$ (resp. $X_{\infty}$) connecting $x$ and $y$. It is known that $(X_{0},d_{0})$ and $(X_{\infty},d_{\infty})$ are complete separable geodesic metric spaces, Moreover, $d_{0}$ and ${d_{\mathbb{R}^{2}}}|_{X_{0}}$, $d_{\infty}$ and ${d_{\mathbb{R}^{2}}}|_{X_{\infty}}$ are bi-Lipschitz equivalent [Bar13].
  2. (2)
    Let
    $\alpha_{{\mathrm{SC}}}:=\frac{\log 8}{\log 3}\text{ and }\beta_{\mathrm{SC}}:=\frac{\log(8\rho)}{\log 3}$
    where $\rho$ is the ‘resistance scaling factor’ of standard Sierpiński carpet so that $\beta_{\mathrm{SC}}$ is its walk dimension [BB92].
  3. (3)
    Let $m_{0}$ be the Lebesgue measure on $\mathbb{R}^{2}$. We define $V_{0}({r})=r^{2}\wedge r^{\alpha_{{\mathrm{SC}}}}$, $r>0$. By [BB00, (2.2)],
    $m_{0}(B_{0}(x,r))\asymp V(r),\ x\in X_{0},\ r>0.$ (8.3)
    where $B_{0}(x,r)=\{y\in X_{0}:d_{0}(x,y)<r\}$.
  4. (4)
    It is known from [BB00, Theorem 5.3] that the heat kernel of the reflected Dirichlet form $(\mathcal{E}_{0},\mathcal{F}_{0})$ on $L^{2}(X_{0},m_{0})$ defined by
    $\mathcal{E}_{0}(u,v):=\int_{\mathcal{P}}\nabla u\nabla v\mathop{}\!\mathrm{d}m_{0},\ u,v\in\mathcal{F}_{0}:=W^{1,2}(X_{0})$
    satisfies $\operatorname{HKE}(\Psi_{0})$, where $\Psi_{0}(r)=r^{2}\vee r^{\beta_{\mathrm{SC}}}$, $r>0$.
  5. (5)
    We define a sequence of metric measure spaces $(X_{n},d_{n},m_{n})$, $n\geq 1$, by
    $X_{n}:=X_{0},\ d_{n}:=l^{-n}d_{0}\ \text{and}\ m_{n}:=l^{-\alpha_{\mathrm{SC}}n}m_{0}.$ (8.4)
    By (8.3),
    $m_{n}(B_{n}(x,r))\asymp V_{n}(r)\ \text{for all $x\in X_{n}=X_{0}$, all $r>0$ and all $n\geq 0$}.$ (8.5)
    where $B_{n}(x,r):=\{y\in X_{n}:d_{n}(x,y)<r\}$ and
    $V_{n}(r):=l^{-\alpha_{\mathrm{SC}}n}V_{0}(l^{n}r).$ (8.6)
    We define a sequence of Dirichlet forms $(\mathcal{E}_{n},\mathcal{F}_{n})$ on $L^{2}(X_{n},m_{n})$ by
    $\mathcal{E}_{n}(u,v):=l^{(\beta_{\mathrm{SC}}-\alpha_{SC})n}\mathcal{E}_{0}(u,v),\ u,v\in\mathcal{F}_{n}:=\mathcal{F}_{0}.$ (8.7)
    It is easy to verify that the heat kernel $p^{(n)}(t,x,y)$ of $(\mathcal{E}_{n},\mathcal{F}_{n})$ with respect to $m_{n}$ is related to the heat kernel $p(t,x,y)$ of $(\mathcal{E}_{0},\mathcal{F}_{0})$ with respect to $m_{0}$ in the following way:
    $p^{(n)}(t,x,y)=l^{\alpha_{\mathrm{SC}}n}p(l^{\beta_{\mathrm{SC}}n}t,x,y),\ x,y\in X_{n}=X_{0},\ t>0\ \text{and}\ n\geq 0.$ (8.8)
    Let
    $\Psi_{n}(r):=l^{-\beta_{\mathrm{SC}}n}\Psi_{0}(l^{n}r),\ r\geq 0,\ n\geq 1.$ (8.9)
    so we have
    $\Psi_{n}^{-1}(t)$ $=l^{-n}\Psi_{0}^{-1}(l^{\beta_{\mathrm{SC}}n}t),$ (8.10)
    $\Phi_{n}(s)=\Phi_{\Psi_{n}}(s)$ $=l^{\beta_{\mathrm{SC}}n}\Phi_{0}(l^{(1-\beta_{\mathrm{SC}})n}s).$ (8.11)
    Combining (8.5), (8.6), (8.8), (8.9), (8.10) and (8.11) we see that the metric measure Dirichlet space $(X_{n},d_{n},m_{n},\mathcal{E}_{n},\mathcal{F}_{n})$ satisfies $\operatorname{HKE}(\Psi_{n})$ with constants independent of $n$. It is easy to verify that conditions (A1)-(A5) holds with $\beta=2$, $\beta^{\prime}=\beta_{\mathrm{SC}}$, $V_{l}(r)=V_{0}(r)$, $V_{u}(r)=V_{\infty}(r)=r^{\alpha_{\mathrm{SC}}}$ and $\Psi_{\infty}(r)=r^{\beta_{\mathrm{SC}}}$.
  6. (6)
    We will prove that $(X_{n},d_{n},q_{0})$ pointed Gromov-Hausdorff converge to $(X_{\infty},d_{\infty},q_{0})$. Define
    $\widetilde{X}_{n}:=l^{-n}X_{0}=\{l^{-n}x:x\in X_{0}\}\subset X_{0},\ n\geq 0,$
    which is called the unbounded scaled Sierpiński carpet in [BB00], so that $\bigcap_{n=1}^{\infty}X_{n}=X_{\infty}$, see Figure 8.1. Define a metric on $\widetilde{X}_{n}$:
    $\widetilde{d}_{n}(x,y):=l^{-n}d_{0}(l^{n}x,l^{n}y),\ \forall x,y\in\widetilde{X}_{n}.$
    Hence the map $x\mapsto l^{-n}x$ is a bijective isometry between $(X_{n},d_{n})$ and $(\widetilde{X}_{n},\widetilde{d}_{n})$. It is easy to see that $\widetilde{d}_{n}$ is a geodesic metric on $\widetilde{X}_{n}$ and $\widetilde{d}_{n}(x,y)$ is the length (with respect to $d_{\mathbb{R}^{2}}$) of shortest path of in $\widetilde{X}_{n}$ connecting $x$ and $y$. It suffices to show that $(\widetilde{X}_{n},\widetilde{d}_{n},q_{0})$ pointed Gromov-Hausdorff converge to $(X_{\infty},d_{\infty},q_{0})$, in other words, the standard planer Sierpiński carpet $X_{\infty}$ is an asymptotic cone [BBI01, Definition 8.2.7] of the pre-carpet $X_{0}$ at $q_{0}$. By joining geodesic in $\widetilde{d}_{n}$ for each $n$ and using compactness, we see that
    $\sup_{x,y\in X_{\infty}}|d_{\infty}(x,y)-\widetilde{d}_{n}(x,y)|\lesssim l^{-n}.$ (8.12)
    In fact, it is obvious that $\widetilde{d}_{n}(x,y)\leq d_{\infty}(x,y)$ for all $x,y\in X_{\infty}$ since $X_{\infty}\subset\widetilde{X}_{n}$; on the other hand, if we define $H_{n}$ to be the union of boundaries of the closed squares of side $l^{-n}$ with corners in $l^{-n}\mathbb{Z}^{2}$ that are contained in $\widetilde{X}_{n}$, then $H_{n}\subset X_{\infty}$ for all $n\geq 1$. For any $x\in X_{\infty}$, let $v_{n}(x)\in H_{n}$ denote the lower left hand corner of one of the squares containing $x$ so that $d_{\infty}(x,v_{n}(x))\leq Cl^{-n}$ for some $C$ by [BB92, (7.4)]. For $x,y\in X_{\infty}$, $d_{\infty}(v_{n}(x),v_{n}(y))$ is less than the length of the shortest path in $H_{n}$ connecting $v_{n}(x)$ and $v_{n}(x)$, which equals to $\widetilde{d}_{n}(v_{n}(x),v_{n}(y))$, since the the shortest path in $\widetilde{X}_{n}$ can be chosen to be piecewise segments that are piecewise parallel to the shortest path in $H_{n}$. Therefore
    $d_{\infty}(x,y)$ $\leq d_{\infty}(x,v_{n}(x))+d_{\infty}(v_{n}(x),v_{n}(y))+d_{\infty}(v_{n}(y),y)$ (8.13)
    $\leq\widetilde{d}_{n}(v_{n}(x),v_{n}(y))+2Cl^{-n}$ (8.14)
    $\leq\widetilde{d}_{n}(x,y)+4Cl^{-n}$ (8.15)
    which gives (8.12).
    It is easy to calculate that the Hausdorff distance between $\widetilde{X}_{n}$ and $X_{\infty}$ is $d_{\mathrm{H}}(\widetilde{X}_{n},X_{\infty})\lesssim l^{-n}$. For each $x\in\widetilde{X}_{n}$, we choose $f_{n}(x)\in X_{\infty}$ such that $d_{n}(x,f_{n}(x))=\min_{y\in X_{\infty}}d_{n}(x,y)$. Obviously, $f_{n}(q_{0})=q_{0}$ for all $n\geq 1$. For $x,y\in\widetilde{X}_{n}$, we have by triangle inequality that
    |$\phantom{\leq\ }\absolutevalue{d_{\infty}(f_{n}(x),f_{n}(y))-\widetilde{d}_{n}(x,y)}$ | (8.16)
    |$\leq\absolutevalue{d_{\infty}(f_{n}(x),f_{n}(y))-\widetilde{d}_{n}(f_{n}(x),f_{n}(y))}+\absolutevalue{\widetilde{d}_{n}(f_{n}(x),x)-\widetilde{d}_{n}(f_{n}(y),y)}\lesssim l^{-n}$ (8.17)
    Moreover, $B_{\infty}(q_{0},r)\subset N_{Cl^{-n}}(f_{n}(B_{n}(q_{0},r)))$ for $C$ large enough. Therefore by Definition 2.1, $(\widetilde{X}_{n},\widetilde{d}_{n},q_{0})$ pointed Gromov-Hausdorff converge to $(X_{\infty},d_{\infty},q_{0})$.
    Figure 8.1: Part of $(\widetilde{X}_{n},\widetilde{d}_{n})$, $1\leq n\leq 3$.
  7. (7)
    We can now apply Theorem 1.2 and obtain a conservative, regular and strongly local Dirichlet form on the unbounded Sierpiński carpet $X_{\infty}$ satisfying $\operatorname{HKE}(r^{\beta_{\mathrm{SC}}})$ and is the Mosco limit of a subsequence of $\{\mathcal{E}_{n}\}$. Compared to [BB89], this gives another construction of diffusion process on Sierpiński carpet. This argument can be adapted to construct diffusion on generalized Sierpiński carpet, as long as the heat kernel bounds on reflected Dirichlet spaces are known.
Example 8.2.
In this example, we re-construct a diffusion process on unbounded Sierpiński gasket from the diffusion process on the unbounded Sierpiński gasket cable system.
Let $p_{0}:=(0,0)$, $p_{1}:=(1,0)$, $p_{3}:=(1/2,\sqrt{3}/2)$, $V_{0}=\{p_{k}\}_{k=1}^{3}$ and let $G_{0}$ be the triangle with vertices $V_{0}$. Define the similitudes $g_{k}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ by $g_{k}(x):=x/2+p_{k}$, $1\leq k\leq 3$. Let $V_{n+1}:=\bigcup_{k=1}^{3}g_{k}(V_{n})$, $G_{n}:=\bigcup_{k=1}^{3}g_{k}(G_{n})$ for $n\geq 1$. Define $G:=\bigcap_{n=0}^{\infty}G_{n}$, $W_{\infty}:=\bigcup_{n=0}^{\infty}2^{n}G$. We call $W_{\infty}$ the unbounded Sierpiński gasket. Let $d$ be the standard Euclidean distance on $\mathbb{R}^{2}$ and $d_{\infty}$ be the geodesic distance on $W_{\infty}$ with respect to $d$. Define $V_{*}:=\bigcup_{n=0}^{\infty}2^{n}V_{n}$ and $E:=\{(x,y)\in V_{*}\times V_{*}:d(x,y)=1\}$. Then $(V_{*},E)$ is an infinite, locally finite, connected graph. By replacing each edge in $E$ by an isometric copy of the line segment $[0,1]$ (called cable) and gluing them in an obvious way at the vertices, we obtain an unbounded connected closed set $W_{0}\subset\mathbb{R}^{2}$, called the corresponding cable system of $(V_{*},E)$. We define a metric $d_{0}$ on $W_{0}$ by using Euclidean distance on each cable and extend $d_{0}$ to a metric on $W_{0}$ which agrees with the graph distance on $V_{*}$. It is easy to see that $d_{0}$ is a geodesic distance. Let $m_{0}$ be the Hausdorff measure on $(W_{0},d_{0})$, which coincides with the one-dimensional Lebesgue measure on each cable. Let
$\alpha_{\mathrm{SG}}:=\frac{\log 3}{\log 2}\text{ and }\beta_{\mathrm{SG}}:=\frac{\log 5}{\log 2}.$
We define $V_{0}(r):=r\vee r^{\alpha_{\mathrm{SG}}}$ and $\Psi_{0}(r):=r^{2}\vee r^{\beta_{\mathrm{SG}}}$. It is known that
$m_{0}(B_{0}(x,r))\asymp V_{0}(r),\ r\geq 0,$
see [DRY23, p.15562] and [Bar98, Lemma 2.1]. In [BB04] (also [BM18, Section 6]) a conservative, strongly local regular Dirichlet form $(\mathcal{E}_{0},\mathcal{F}_{0})$ on $L^{2}(W_{0},m_{0})$ is defined, and the corresponded continuous Markov process is called the cable process. It is known from [Bar03, Theorem 2.4] and [BB04] (see also [DRY23, p.15563]) that the metric measure Dirichlet space $(W_{0},d_{0},m_{0},\mathcal{E}_{0},\mathcal{F}_{0})$ satifies $\operatorname{HKE}(\Psi_{0})$. We define a sequence of metric measure Dirichlet spaces $(W_{n},d_{n},m_{n},\mathcal{E}_{n},\mathcal{F}_{n})$, $n\geq 1$, by
$W_{n}:=W_{0},\ d_{n}:=2^{-n}d_{0}\ \text{and}\ m_{n}:=2^{-\alpha_{\mathrm{SG}}n}m_{0}$ (8.18)
and
$\mathcal{E}_{n}(u,v):=2^{(\beta_{\mathrm{SG}}-\alpha_{SG})n}\mathcal{E}_{0}(u,v),\ u,v\in\mathcal{F}_{n}:=\mathcal{F}_{0}.$ (8.19)
Similar to Example 8.1-(5), we see that $(W_{n},d_{n},m_{n},\mathcal{E}_{n},\mathcal{F}_{n})$ statifies $\operatorname{HKE}(\Psi_{n})$ with uniform constants, where
$\Psi_{n}(r):=2^{-\beta_{\mathrm{SG}}n}\Psi_{0}(2^{n}r),\ r\geq 0,\ n\geq 1,$ (8.20)
and conditions (A1)-(A5) holds with $\beta=2$, $\beta^{\prime}=\beta_{\mathrm{SG}}$, $V_{l}(r)=V_{0}(r)$, $V_{u}(r)=V_{\infty}(r)=r^{\alpha_{\mathrm{SG}}}$ and $\Psi_{\infty}(r)=r^{\beta_{\mathrm{SG}}}$. Similar to Example 8.1-(6), it is easy to verify that $(W_{n},d_{n},p_{0})$ pointed Gromov-Hausdorff converge to $(W_{\infty},d_{\infty},p_{0})$. By Theorem 1.2, we obtain a conservative, regular and strongly local Dirichlet form on the unbounded Sierpiński gasket $W_{\infty}$ satisfying $\operatorname{HKE}(r^{\beta_{\mathrm{SG}}})$ and is the Mosco limit of a subsequence of $\{\mathcal{E}_{n}\}$.
Figure 8.2: Part of the cable systems.
Example 8.3.
Barlow in [Bar04] shows that, the existence of an infinite locally finite weighted graph that satisfying $\mathrm{V}_{\alpha}$, $\mathrm{EHI}$ and $\mathrm{E}_{\beta}$, which is equivalent to the random walk satisfying the following type heat kernel estimates: for some $C_{1},c_{2}\geq 1$ and all $n\geq 1$,
$p_{n}(x,y)$ $\leq\frac{C_{1}}{n^{\alpha/\beta}}\exp\left(-c_{2}\left(\frac{d(x,y)^{\beta}}{n}\right)^{1/(\beta-1)}\right)$ (8.21)
and $\text{and \quad}p_{n}(x,y)+p_{n+1}(x,y)$ $\geq\frac{C^{-1}_{1}}{n^{\alpha/\beta}}\exp\left(-c^{-1}_{2}\left(\frac{d(x,y)^{\beta}}{n}\right)^{1/(\beta-1)}\right),$
is equivalent to
$\alpha\geq 1\text{ and }2\leq\beta\leq\alpha+1.$ (8.22)
In a recent paper by Murugan [Mur24], the analogous result on diffusion on metric measure space is proved: the existence of a metric measure Dirichlet space satisfying the full heat kernel bounds $\operatorname{HKE}_{\mathrm{f}}(\Psi)$ with $m(B(x,r))\asymp V(r)$ is equivalent to
[…]
see Remark 1.3-(i) for definition of $\operatorname{HKE}_{\mathrm{f}}(\Psi)$. In the spacial case of $V(r)=r^{\alpha}$ and $\Psi(r)=r^{\beta}$, (8.23) is equivalent to (8.22). In this example, we give an independent proof of a special case that (8.22) implying the existence of such metric measure Dirichlet space, as an application of Theorem 1.2. The method we use here is to take the pointed Gromov-Hausdorff limit of the cable systems of the graphs constructed by Barlow in [Bar04].
Theorem 8.4 ([Mur24, Theorem 2.6]).
For any $\alpha\geq 1\text{ and }2\leq\beta\leq\alpha+1$, there is a metric measure space $(X,d,m)$ and a conservative, regular and strongly local symmetric Dirichlet form on $L^{2}(X,m)$ that satisfies $\operatorname{HKE}_{\mathrm{f}}(\Psi)$ with $\Psi(r)=r^{\beta}$ and $m(B(x,r))\asymp r^{\alpha}$.
Proof.
  1. (1)
    Fix $\alpha$ and $\beta$ such that $\alpha\geq 1\text{ and }2\leq\beta\leq\alpha+1$. By [Bar04] and [GT02, Theorem 3.1], there is an infinite locally finite weighted graph $\Gamma=(V,E,\mu)$ satisfying (8.21). It can be directly checked that $\Gamma$ also satisfies the $p_{0}$-condition stated in [BB04, (1.5)] (or use the heat kernel lower bound in (8.21)). Let us denote the cable system corresponding to $\Gamma$ by $\Gamma^{(c)}$. Similar to Example 8.2, we equip $\Gamma^{(c)}$ a metric $d$ by extending the Euclidean distance on each cable, and let $m$ be the Hausdorff measure on $\Gamma^{(c)}$, so that $(\Gamma^{(c)},d,m)$ is a proper metric measure space. The Dirichlet form corresponding to the cable process on $(\Gamma^{(c)},d,m)$ is denoted by $(\mathcal{E},\mathcal{F})$.
  2. (2)
    We denote $V(r):=r\vee r^{\alpha}$ and $\Psi(r):=r^{2}\vee r^{\beta}$. By [BB04, Lemma 2.1],
    $m(B(x,r))\asymp V(r),\ r\geq 0.$
    The elliptic Harnack inequality $\mathrm{EHI}$ holds on $(\mathcal{E},\mathcal{F})$ by [BB04, Corollary 2.5]. The mean exit time estimate $\mathrm{E}_{\beta}$ on cable process holds by [BB04, Lemma 2.6]. Therefore we can apply [GT12, Theorem 5.15] and conclude that $(\mathcal{E},\mathcal{F})$ on $L^{2}(\Gamma^{(c)},m)$ satisfies $\operatorname{HKE}(\Psi)$. Fix any sequence of positive real numbers $\{\varepsilon_{n}\}_{n=1}^{\infty}$ with $\varepsilon_{1}=1$ and $\varepsilon_{n}\downarrow 0$. We define a sequence of metric measure Dirichlet spaces $(\Gamma^{(c)}_{n},d_{n},m_{n},\mathcal{E}_{n},\mathcal{F}_{n})$, $n\geq 1$, by
    $\Gamma^{(c)}_{n}:=\Gamma^{(c)},\ d_{n}:=\varepsilon_{n}d\ \text{and}\ m_{n}:=\varepsilon_{n}^{\alpha}m$ (8.24)
    and
    $\mathcal{E}_{n}(u,v):=\varepsilon_{n}^{\beta-\alpha}\mathcal{E}(u,v),\ u,v\in\mathcal{F}_{n}:=\mathcal{F}.$ (8.25)
    Define $V_{n}(r):=\varepsilon_{n}^{\alpha}V(\varepsilon_{n}^{-1}r)$ and $\Psi_{n}(r):=\varepsilon_{n}^{\beta}\Psi(\varepsilon_{n}^{-1}r)$ for $r\geq 0$ and $n\geq 1$. Then the metric measure Dirichlet space $(\Gamma^{(c)}_{n},d_{n},m_{n},\mathcal{E}_{n},\mathcal{F}_{n})$ satisfies (A1)-(A5) with $V_{\infty}(r)=r^{\alpha}$, $\Psi_{\infty}(r)=r^{\beta}$.
  3. (3)
    Fix $p\in\Gamma^{(c)}=\Gamma^{(c)}_{n}$. We claim that $(\Gamma^{(c)}_{n},d_{n},p)$ is pointed totally bounded [HKST15, p.321]: there is a function $N:(0,\infty)\times(0,\infty)\to(0,\infty)$ such that, for each $0<r<R$ and $n\geq 1$, the closed ball $B_{n}(p,R)$ in $(\Gamma^{(c)}_{n},d_{n})$ contains an $r$-net of cardinality at most $N(r,R)$. Since in the sense of sets, $B_{n}(p,R)$ is identical to $B_{1}(p,\varepsilon_{n}^{-1}R)$, which, by Zorn’s Lemma, contains an $\varepsilon_{n}^{-1}r$-net of cardinality at most $C(R/r)^{\alpha}$ by the existence of doubling measure $m$ and [Hei01, Exercise 10.17] for $C\geq 1$ independent of $n$. We thus may take $N(r,R)=C(R/r)^{\alpha}$.
  4. (4)
    Since $(\Gamma^{(c)}_{n},d_{n},p)$ are proper length and pointed totally bounded, we conclude by Gromov’s compactness theorem [HKST15, Theorem 11.3.16] that there is a proper length pointed metric space $(\Gamma^{(c)}_{\infty},d_{\infty},p_{\infty})$ such that
    $(\Gamma^{(c)}_{n},d_{n},p)\xrightarrow{\mathrm{p-GH}}(\Gamma^{(c)}_{\infty},d_{\infty},p_{\infty})$
    along some subsequence. Therefore (A6) holds along this subsequence.
  5. (5)
    Now it is time to apply Theorem 1.2 and Remark 1.3-(i), and conclude that there is a measure $m_{\infty}$ on $(\Gamma^{(c)}_{\infty},d_{\infty})$ that is Ahlfors $\alpha$-regular, i.e., $m_{\infty}(B_{\infty}(x,r))\asymp r^{\alpha}$, and there is a conservative, regular and strongly local symmetric Dirichlet form on $L^{2}(\Gamma^{(c)}_{\infty},m_{\infty})$ that satisfies heat kernel bounds $\operatorname{HKE}_{\mathrm{f}}(r^{\beta})$.
Acknowledgments. I am very grateful to Mathav Murugan for introducing me to the problem of constructing a Dirichlet form on limit space, for the suggestion to apply the main result to Example 8.3, and for many helpful discussions regarding this topic. I thank Jiaxin Hu for his helpful comments and interest in this paper.